### v1

# 1]FAIR at Meta

Abstract
[…]
1 Introduction
Over the past few years, Transformer-based architectures (Vaswani et al., 2017) have demonstrated impressive performance in different tasks, including holding conversations at the human level (Shuster et al., 2022; OpenAI, 2022, 2023; Touvron et al., 2023), high-quality image understanding (Caron et al., 2021; Oquab et al., 2024; Assran et al., 2023) and video generation (Singer et al., 2023), multi-modal generation (Girdhar et al., 2023; Radford et al., 2021), and code completion (Roziere et al., 2023; OpenAI, 2021). By training these architectures with a huge amount of data, the resulting models, such as Large Language Models (LLMs), can generalize well in real-world use cases.
Despite these successes, Transformer-based architectures and LLMs still struggle when it comes to solving planning and reasoning tasks. Previous studies (Momennejad et al., 2023; Valmeekam et al., 2023a, b) demonstrate that LLMs fall short in multi-step planning tasks (Valmeekam et al., 2023b) or when performing higher-order reasoning (Momennejad et al., 2023; Fan et al., 2020).
In recent years, various methods have been proposed to improve the performance of Transformers in reasoning and planning tasks. One common and effective approach is to simulate the human thinking process and produce intermediate “thoughts” before outputting a response. Chain-of-Thought (CoT) prompting (Wei et al., 2022) encourages the model to predict the intermediate steps and to “think” step by step. Tree-of-thoughts (ToT) uses a branching strategy and critics to generate different thought paths before picking the best one (Yao et al., 2023). While these techniques are often effective, there are studies showing that in many cases, they may lead to worse performance, for example due to self-enforcing (Huang et al., 2023). Furthermore, techniques effective on one dataset may not work well on others due to changes in the type of reasoning involved (e.g., spatial reasoning vs. mathematical reasoning vs. common-sense reasoning). How to enable Transformers or LLMs to plan, solve multi-step decision making tasks, and perform different types of reasoning still remains elusive and an active area of research.
These methods stand in sharp contrast with traditional symbolic planning and search techniques. While such techniques may not exhibit the language understanding capabilities of LLMs trained on internet-scale datasets, they are well-defined and can solve planning tasks that are far more complex than the tasks state-of-the-art LLMs can solve (Momennejad et al., 2023; Valmeekam et al., 2023a). Additionally, there often exist formal guarantees for solutions computed by these traditional methods, because symbolic planning algorithms often follow a well-defined rule-based search procedure.
Our work
In this paper, we demonstrate how to train Transformers to solve complex planning tasks and present Searchformer, a Transformer model that computes an optimal plan in fewer search steps than symbolic planning algorithms such as $A^{*}$ search (Russell and Norvig, 2021, Chapter 3) in multi-step planning tasks such as maze navigation and solving Sokoban puzzles. To achieve this, we introduce search dynamics bootstrapping, a method that first trains a Transformer model to imitate the search procedure of $A^{*}$ and then fine-tunes the model to find an optimal plan with fewer search steps.
In the first step a Transformer model is trained to imitate $A^{*}$ search. Here, we generate synthetic datasets by running $A^{*}$ search in randomly generated planning task instances. While $A^{*}$ is executed, we log the performed computation and optimal plan into a sequence of words, called tokens. The resulting training dataset then contains an execution trace of $A^{*}$ and encodes information about the search dynamics of $A^{*}$ itself. A Transformer model is then trained to generate these token sequences along with an optimal plan for any given planning task.
In the second step, we show that our Searchformer trained with search-augmented sequences—sequences that include $A^{*}$’s execution trace—can be further improved via expert iteration (Gulcehre et al., 2023), a method that encourages the Transformer to generate fewer search steps while still generating optimal solutions. This procedure leads to a neural planning algorithm that is implicitly encoded in the Transformer’s network weights, and finds an optimal plan with high probability in fewer search steps than standard $A^{*}$ search. Specifically, when solving Sokoban puzzles, our models solve 93.7% of all test tasks while performing search steps that are on average 26.8% shorter than $A^{*}$ search. This paves the way to use Transformers to move beyond traditional symbolic planning algorithms.
To better understand how the used training data and number of model parameters influence performance of the resulting models, we present a set of ablation studies. We train models on both solution-only sequences (token sequences including task description and final plan) and search-augmented sequences (token sequences including task description, search tree dynamics, and final plan) and compare their performance. Despite the fact that search-augmented sequences are thousands of tokens long, search-augmented models (that are trained on search-augmented sequences) can still learn to generate the correct search dynamics and optimal plans with high probability. Further, compared to solution-only models (that are trained on solution-only sequences), search-augmented models generate an optimal plan more often on unseen tasks with ten times fewer training sequences, highlighting the power of including $A^{*}$’s search dynamics into the training process of Transformer models. Since the search dynamics can be included into the training data simply by through logging when generating synthetic training sequences, our study provides a data-efficient way to improve the reasoning power of Transformer models. We also find that solution-only models do not improve with more parameters, showing that the type of training data used plays a critical role in this setting.
2 Related Work
While existing work (Trinh et al., 2024; Ruoss et al., 2024) leverages synthetic datasets to learn strong policies for reasoning, our study is fundamentally different in this regard. We focus on improving the reasoning capability embedded in a Transformer’s weights, while algorithms such as AlphaZero (Silver et al., 2018), MuZero (Schrittwieser et al., 2020) and AlphaGeometry (Trinh et al., 2024) treat general neural network models (including Transformers) as black-boxes, and use existing symbolic planning techniques to improve the neural network itself. For example, Silver et al. (2017) use MCTS as a policy improvement operator to update the neural network’s weights. In contrast, the search dynamics bootstrapping method in this work uses a Transformer model to generalize towards more efficient search patterns and improve the model itself. A planning algorithm is only used to initially train a Transformer model.
ScratchPad (Nye et al., 2021) train neural networks on execution traces of reasoning tasks, Yang et al. (2022), Pallagani et al. (2022), and Ruoss et al. (2024) train models to predict the optimal solution or an optimal action directly. In contrast, we find more efficient LLM-based planning algorithms by predicting the search dynamics of a planning algorithm, which is the execution trace when the solver attempts to find the optimal solution, which is a harder problem than predicting an optimal decision making strategy.
Our work also bears some similarity with neuro-symbolic systems (e.g., Neural Turing Machines (Graves et al., 2014), Neural Programming Architecture (Cai et al., 2017)), which build differentiable architectures to mimic the functionality of existing symbolic systems. However, these methods use dedicated components (e.g., explicit memory components, built-in recursion), while Searchformer uses a general Transformer architecture. By using long contexts and positional embeddings (Chen et al., 2023; Peng et al., 2023) to predict an optimal plan, Searchformer performs well in solving different complex planning tasks. Ultimately, our work is a step towards building more general neural architectures that automatically learn a reasoning and planning mechanism without prior human knowledge.
: wall cell: start cell: goal cell: plan step210210
(a) Maze navigation task
Promptbosstart 0 2goal 1 0wall 1 2wall 2 0eosResponsebosplan 0 2plan 0 1plan 1 1plan 1 0eos
(b) Tokenization of a planning task and its solution
[…]
(c) $A^{*}$’s execution when solving a planning task is logged into an execution trace
Figure 1: Expressing a planning task in token sequences. 1(a): A $3\times 3$ maze navigation task where an agent must navigate from the start cell to the goal cell without entering a wall cell. For each transition, a cost of one is given and the optimal plan is the shortest path from start to goal. 1(b): The $3\times 3$ maze navigation task is expressed as a prompt token sequence (left panel) and the optimal plan is expressed as a response token sequence (right panel). The start and end of a response sequence is indicated by a beginning-of-sequence token, bos, and an end-of-sequence tokens, eos. The numbers indicate $x,y$ coordinates. 1(c): The search dynamics of the $A^{*}$ algorithm (left panel) is logged into an execution trace represented by a token sequence (right panel). For maze tasks, coordinates are encoded with integer tokens. The next two numbers costs associated with that node (letter “c” is added to differentiate them from coordinate numbers). Each token is mapped to a unique non-negative integer and this integer is processed in the neural network.
Using Transformer architectures to solve complex sequential decision making has been studied in prior work in a reinforcement learning (RL) setting (Chen et al., 2021; Janner et al., 2021; Laskin et al., 2023). Decision Transformer (Chen et al., 2021) predicts the next optimal action given a trajectory segment that is annotated with total return values, while Trajectory Transformer (Janner et al., 2021) focuses on learning a world model that predicts future states, actions, and rewards. Similarly, in Algorithm Distillation (Laskin et al., 2023) a Transformer architecture is trained to predict the next action given a trajectory segment of trial and error interactions as context. In contrast, we incorporate how a planning algorithm arrives at a final result into a Transformer models. MCTSNet (Guez et al., 2018) also attempts to learn the search procedure itself, but still hard-codes the MCTS search procedure into neural network, which leads to quadratic backpropagation overhead and can only deal witih up to 500 rollouts, while our approach can deal with much longer search execution trace. We demonstrate that Transformers can not only imitate a symbolic planning algorithm but also improve past the algorithm it was initially trained on via fine-tuning.
3 Problem Setup
Figure 2: Example Sokoban puzzle [^2] This level image was composed using image icons from https://github.com/morenod/sokoban (accessed 2023-11-21).. The prompt, $A^{*}$ execution trace, and optimal plan for this task is illustrated in Figure 10 in Appendix A.3. For Sokoban, we only use a non-deterministic $A^{*}$ implementation.
We consider two domains: maze navigation tasks (Figures 1(a)) and solving Sokoban puzzles (Figure 2). In maze navigation, the goal is to find the shortest path in an $n$-by-$n$ maze between a start location and a goal location without passing through wall cells. Figure 2 depicts a Sokoban puzzle where the worker can move up, down, left, or right to push each box onto a dock (crossed grid cells). The worker cannot pull a box away from a wall cell. To solve the puzzle, each box must be moved onto a dock in as few steps as possible. An incorrect move may immediately lead to a dead end and thus careful planning is needed to solve the task. Furthermore, each state in a Sokoban puzzle consists of every possible combination of a box and worker position. Consequently, state spaces are large making a Sokoban puzzle computationally more difficult to solve for algorithms such as $A^{*}$ search.
3.1 Generating execution traces of $A^{*}$ search.
The $A^{*}$ algorithm computes an optimal plan by manipulating two sets of nodes:
  • A frontier set containing the current search frontiers;
  • A closed set containing all searched nodes.
In the maze example in Figure 1(a), each node corresponds to an empty (non-wall) grid cell. For each node, the algorithm computes a heuristic value and a cost from start value. At any given iteration, which node is searched next is determined by the content of the frontier and closed sets as well as the heuristic and cost from start values (Figure 1(c), left panel). The $A^{*}$’s execution trace is collected by tracking all insertion operations into the frontier and closed set along with heuristic and cost from start values. The right panel in Figure 1(c) illustrates the resulting trace for the maze example shown in Figure 1(b). Each row corresponds either to an insertion of a node into the frontier, indicated by a create token, or to moving a node into the closed set, indicated by a close token. Each node is represented by its $(x,y)$ position in the maze as well as the two cost tokens. The resulting plan is then appended to this trace. This trace is constructed such that given any prefix of this trace the next token can be predicted correctly. For the maze datasets, $A^{*}$ uses the Manhattan distance to the goal location as a heuristic. In Sokoban, $A^{*}$ first matches every box to the closest dock and then computes the sum of all Manhattan distances between each box and dock pair.
For each experiment, we generate two token sequence variants, as illustrated in Figure 1:
  • Search-augmented sequence. This variant contains sequences of the format <prompt><trace><plan>, where the <trace> part encodes $A^{*}$ execution trace and <plan> encodes the optimal plan.
  • Solution-only sequence. This variant omits $A^{*}$’s exeuction trace and only contains sequences of the format <prompt><plan>.
For example, the $3\times 3$ maze navigation task (Figure 1(a)) can be mapped into a prompt and plan sequence (Figure 1(b)). The <trace> part is constructed by running $A^{*}$ on each planning task and logging the performed computations into a token sequence (Figure 1(c)). While the optimal plan in both cases can be fairly short, the corresponding execution traces can be quite long, in particular when $A^{*}$’s heuristic underestimates the cost to reach the goal state and often misleads the search to explore in-feasible paths.
As illustrated in Figure 1, all token sequences are synthetically generated and use a special vocabulary. Furthermore, every model is trained from scratch—no natural language dataset is used for training. Consequently, the resulting models are specifically trained to only predict sequences that outline optimal plans for a set of different planning tasks. After training, the model’s output is parsed and evaluated if it contains a correct and optimal solution plan.
3.2 Training a Transformer model
To train a Transformer model, we randomly generate a set of planning tasks together with their corresponding token sequences. In each dataset, each task is unique and the test set is constructed such that it does not contain any duplicate of the training set. For example, a maze task that occurs in the training dataset does not occur in the test dataset and no task is repeated. With this experiment design, we hope to gain insight into how Transformers can be used to solve planning tasks.
By including intermediate computation steps, the Transformer model is trained to effectively imitate the computation performed by the $A^{*}$ algorithm. This approach to training a neural network is similar to Procedure Cloning (Yang et al., 2022) where a neural network is not only presented with input-output pairs (in our case task prompts and optimal plans), but also with examples of how to “think” about a particular task
In our experiments, we train an adaptation of the encoder-decoder T5 architecture (Raffel et al., 2020) and use Rotary Position Embeddings (RoPE) (Su et al., 2023). More details and hyper-parameters about the models we use can be found in Appendix A.2. The encoder processes the <prompt> part of a training sequence, and the decoder processes either a <trace><plan> formatted sequence to train a search-augmented model or only a <plan> formatted sequence when training a solution-only model. Depending on the model variant, each network is trained to maximize the cross-entropy between the distribution of the decoder generations and the distribution of sampling a corresponding sequence from the training dataset. Gradient updates are computed using teacher forcing. Appendix A.1 describes our optimization setup in more detail.
3.3 Moving past algorithm imitation via search dynamics bootstrapping
To reduce the number of search steps generated by a search-augmented model during inference, we implement a method to shift the distribution with which the decoder generates execution traces. First, a search-augmented model is trained to imitate the search dynamic of our non-deterministic $A^{*}$ implementation. This non-deterministic $A^{*}$ implementation breaks cost ties randomly and randomizes the order with which child nodes are expanded. By doing so, we do not decrease the efficiency of $A^{*}$ search itself and merely change the order with which different nodes are searched while still respecting $A^{*}$’s heuristic and cost calculations.
Using a non-deterministic implementation randomizes the sequence generation process and induces additional variance into the length of each execution trace. The resulting search-augmented model will then approximate the probability distribution with which the training sequences were generated.
Once a model is trained to imitate the search dynamics of non-deterministic $A^{*}$ search, it is used to generate a new training dataset consisting of shorter token sequences. This new dataset is constructed by using the trained search-augmented model to generate multiple token sequences for each training prompt. In this step, we only use the training dataset for bootstrapping and not the test dataset. Each generated sequence is parsed and checked if it ends in an optimal plan. If this is the case and the sequence is also shorter than the corresponding sequence contained in the original training dataset, then this shortened sequence is included in the new short sequence training dataset. If the generated sequence does not end in an optimal plan or is longer, then the sequence from the original training dataset is re-used.
Subsequently, the search-augmented model is fine-tuned on the new short sequence training dataset. To distinguish from the search-augmented model trained on the original dataset, we call this new model Searchformer. This procedure can then be repeated by using the resulting fine-tuned model to generate the next even shorter sequence dataset and then using this dataset to further fine-tune the Searchformer model. In Section 4.4 we demonstrate that this procedure does in fact reduce the number of search steps performed during inference while further improving performance. At this point, the Searchformer model no longer imitates $A^{*}$ search and has instead discovered a new way of solving a planning problem using fewer search steps.
4 Experiments
In our experiments, we use a deterministic and a non-deterministic variant of $A^{*}$ search to generate each sequence dataset.
Deterministic $A^{*}$ dataset
In the first case, sequences are generated by executing $A^{*}$ in a deterministic fashion (by ordering child nodes and breaking equal cost ties in the frontier set deterministically). Consequently, given a task prompt, the optimal plan and $A^{*}$ execution trace is unique. A Transformer model trained on this data only has to learn which token to predict next. Evaluating such a model is simple, because the generated token sequences need to exactly match the token sequences generated by $A^{*}$.
Non-deterministic $A^{*}$ dataset
Figure 3: Training Sequence Length Comparison. The left panel plots the length of the solution-only sequences and the right panel plots the length of the search-augmented sequences, excluding the start and end of sequence tokens bos and eos. The whiskers indicate the range of all sequence lengths and the box plots indicate the 25%, 50%, and 75% quantiles. Because we focus on planning in complex sequential decision making tasks, the token sequences are multiple orders of magnitude longer than usual token sequences used to train LLMs—especially when $A^{*}$ execution traces are included in the responses. For example, fine-tuning of the Llama 2 model on human preference data is performed with sequences that are on average 600 tokens long (Touvron et al., 2023).
In practice, there can be multiple plans that solve the same task optimally and a deterministic implementation of $A^{*}$ search will only compute one optimal solution. To diversify the dataset and include different search sequences, we generate sequences by executing $A^{*}$ search in a non-deterministic fashion: in each run, we randomly shuffle the order with which child nodes are expanded and break equal cost ties randomly. Consequently, the generated <trace><plan>-formatted sequences vary between different executions. The resulting plans may differ as well but are still optimal. This modification does not decrease the efficiency of $A^{*}$ search, but merely forces our implementation to use different search sequences that still correspond to executing the pseudo-code listed in Figure 1(c). Other randomization strategies—which do influence $A^{*}$’s efficiency—are also possible (e.g., randomizing cost values at each step) and are left for future works.
Statistics
Figure 3 presents an overview of the token sequence length used for each model type for the Sokoban dataset and each maze datasets. The length of the generated $A^{*}$ execution traces grows with the grid size of each maze. In Sokoban, the task with the largest state space, execution traces are the longest. However, the solution plans for Sokoban are amongst the shortest.
4.1 Evaluation Criteria
For each task, one model may generate a search sequence ending either in an optimal plan, a correct but sub-optimal plan, or an invalid plan. For a plan to be valid, we not only check whether the final solution sequence is feasible, but also verify the output format. The search-augmented models must generate a token sequence of the format <trace><plan>, while solution-only models must generate a token sequence of the format <plan> only. Generated sequences violating such constraints—including sequences that do not predict an end-of-sequence token eos—are omitted.
4.1.1 Measuring plan optimality
We evaluate whether the output plan is optimal using one of three criteria:
  • Exact-match criterion. For each task, if the generated sequence from a trained model matches the output of deterministic $A^{*}$ exactly, it is labelled as correct, otherwise labelled as incorrect. This is only used to evaluate supervised cloning of deterministic $A^{*}$.
  • Any-optimal-64 criterion. For each task, we sample 64 responses from a trained model. They are parsed and then evaluated to test if the final plan is feasible and optimal, regardless of the generated <trace> part. If any of the 64 plans is feasible and optimal, then the task is labelled as correct.
  • SWC score. To further measure the sub-optimalness of the resulting plans, we also report the Success Weighted by Cost (SWC) (Wu et al., 2019), a statistic that factors in how close the cost $l_{i}$ of the best predicted correct plan (over 64 trials) is to the optimal plan cost $l_{i}^{*}$, averaged across all $n$ test tasks and weighted by a binary variable $c_{i}\in\{0,1\}$:
    $\text{SWC}:=\frac{1}{n}\sum_{i=1}^{n}c_{i}\frac{l_{i}^{*}}{\max\{l_{i},l_{i}^{*}\}}.$
    When computing the SWC score, the binary variable $c_{i}$ is set to one if a correct plan was found and zero otherwise. This SWC score lies between zero and one. If all generated sequences end in an optimal plan, then this value is one.
4.1.2 Measuring search dynamics length
Given the plan is correct or optimal, we evaluate the length of sequence dynamics in terms of number of tokens using one of the two criteria:
  • Average-on-optimal length. For each task, we sample 64 responses from a trained model and compute averaged search dynamics length for sequences that lead to an optimal plan.
  • ILR score. To further measure how much improvement of the model-generated search dynamics against $A^{*}$ planning, we report the Improved Length Ratio of Search Dynamics (ILR) score. Specifically, for each test task $i$, we compute the ratio between the length $t_{i}$ of the shortest generated search dynamics sequence and the $A^{*}$ token sequence length $t_{i}^{*}$. We then average across all test tasks while only including ratios for tasks for which either an optimal or a correct (and potentially sub-optimal) plan was found, as specified by the binary variable $c_{i}$. The corresponding measures are thus called ILR-on-optimal and ILR-on-solved. The ILR is defined as
    $\text{ILR}:=\frac{1}{n}\sum_{i=1}^{n}c_{i}\frac{t_{i}^{*}}{t_{i}}.$
    The ILR measure can take non-negative values and values above one indicate that the model generates shorter search dynamics than the $A^{*}$ reference. Consequently, if the numbers lie significantly above one, then the model has found a more efficient way to search a task’s state space to compute an optimal plan.
Unless indicated otherwise, each experiment is repeated five times and each figure plots averages across all repeats. Error bars are used to indicate the Standard Error of Measurement (SEM).
4.2 Maze Navigation
(a) Deterministic Prediction
(b) Non-deterministic Prediction
Figure 4: Predicting intermediate computation steps leads to robust performance in the low data regime. Each model was trained to predict optimal plans for $30\times 30$ maze navigation tasks. The decoder of the search-augmented models were trained on sequences of the format <trace><plan> whereas the decoder of the solution-only models were trained on sequence of the format <plan>. For each model, the number of free parameters (indicated in millions of parameters with “15M”, “46M”, and “175M”) is varied by changing the hyper-parameters of the encoder-decoder architecture. Appendix A.2 describes the used architecture and hyper-parameters in more detail.
In the first experiment set, we train a set of encoder-decoder Transformer models to predict optimal plans for $30\times 30$ mazes. We vary the training dataset size and model size (the number of optimized parameters) between different training runs and evaluate each model on the test tasks generated using the same hyper-parameters (e.g., maze size).
Evaluation with the exact-match criterion
Figure 4(a) plots for how many test tasks a correct response was generated. Both solution-only and search-augmented models are trained on the deterministic $A^{*}$ dataset and follow the exact-match criterion. One can observe that the solution-only model is outperformed by most search-augmented models. Only for large enough training datasets, the solution-only model matches the performance of the worst search-augmented model. In the low training data regime (100,000 training sequences and less), performance of the solution-only model degrades significantly, while the performance of each search-augmented model stays relatively high.
This result is surprising, because the search-augmented models must generate a sequence of the format <trace><plan>, which is on average more than 10 times longer than the corresponding sequence of the format <plan> generated by the solution-only models. Furthermore, even the smallest search-augmented model with 15 million (15M) parameters significantly outperforms the much larger solution-only model with 175 million (175M) parameters.
Evaluation with the any-optimal-64 criterion
When trained on non-deterministic $A^{*}$ data, there could be multiple different optimal paths for one task, and in practice one would only use the model to compute an plan, the any-optimal-64 metric is more appropriate. Note that using the any-optimal-64 criterion results in higher absolute numbers compared to exact-match, because to score one point under this criterion only one generated response needs to be correct.
Figure 5: Solution-only model performance. Performance of the solution-only models is primarily influenced by the number of training sequences. Increasing a model’s size does not always improve performance.
Figure 6: Performance across different task difficulties. Increasing task difficulty from left to right degrades performance of the solution-only models significantly. Here, each model was trained on 50,000 sequences.
Figure 4(b) plots the any-optimal-64 criterion on the test set as a function of the number of non-deterministic $A^{*}$ training sequences. We can observe a similar pattern: even the smallest search-augmented models outperform solution-only model, especially for small training set. Moreover, we found that model size only impacts the performance of each of the search-augmented models when using very small training datasets (50,000 training sequences). For larger training dataset sizes no significant difference is found. Increasing the number of parameters of the solution-only models does not significantly improve their performance in the low data regime (Figure 6).
Performance across different task difficulty levels
Figure 6 illustrates how a task’s difficulty influences the performance of each model. Here, we focus again on the dataset generated by non-deterministic $A^{*}$, and consider the number of correctly solved test tasks as a function of maze size. The larger the maze, the larger the task’s state space and the more computation is required to find an optimal solution plan. While the solution-only model reaches close to the performance of search-augmented model for the small $10\times 10$ mazes, its performance drops rapidly as the task becomes more challenging. In contrast, the search-augmented models maintain their high accuracy as the task becomes harder, even for its smallest model size. Appendix A.5 presents A full comparison across all maze sizes.
Figure 7: Performance comparison with the Sokoban dataset. Percentage of correctly answered test prompts for each model size. All models see 100,000 training sequences. The plot averages percentage values across three different seeds and evaluates each model on the same 200 test tasks. Test tasks are not included in the training dataset.
Overall, while the solution-only models learn to predict an optimal plan if the used training dataset is large and diverse enough, search-augmented models perform significantly better in the low data regime and scale better to more difficult tasks. The search-augmented models reach higher performance because they can perform on-demand computation during inference. More specifically, the search-augmented models can control the search dynamics for a grounded reasoning chain that leads to an optimal plan, while the solution-only models have to infer direct correlations between a task description and its optimal plan through supervised learning where many of such correlations can be spurious and unreliable in the test time.
4.3 Solving Sokoban puzzles
To test if similar results can be obtained on a different and more complex task with a different tokenization pattern, we generate a planning dataset for Sokoban puzzles. Here, $A^{*}$’s execution traces are generated non-deterministically and expressed as a token sequence as outlined in Figure 10.
Figure 7 plots how often each model generated a correct optimal plan for each test task. As before, by training on execution traces, the search-augmented models outperform the solution-only models. Even increasing the parameterization of a solution-only model to 747 million parameters only leads to a marginal performance improvement. On average, this 747 million parameter solution-only model is still outperformed slightly by a smaller 175 million parameter search-augmented model. This experiment confirms the findings found for mazes on a dataset with more complex planning tasks that use a different tokenization method.
4.4 Searchformer: Improving search dynamics via bootstrapping
In this last experiment, we investigate how the search-augmented model can be iteratively improved to compute an optimal plan while predicting fewer search steps. Here, our goal is to shorten the length of the search trace while still producing an optimal solution.
We start out with the smallest search-augmented model trained on the non-deterministic $A^{*}$ Sokoban dataset and use it to generate a new shorter sequence training dataset as outlined in Sec. 3.3. For each Sokoban puzzle in the training data, we generated 32 different <trace><plan> sequences by sampling tokens from the Transformer’s output distribution and include the shortest generation (measured in tokens) if it contains an optimal plan.
Subsequently, we fine-tune the search-augmented model on this newly created training data (by running an additional 10,000 training steps) to obtain the first Searchformer model. Using this Searchformer model, we subsequently generate another short sequence dataset and repeat the fine-tuning procedure. This newly generated dataset is then used to further fine-tune the previously obtained Searchformer model for another 10,000 training steps. In total, we repeat this improvement step three times.
(a) Comparison of distribution of search dynamics length between $A^{*}$ search and our models. Each test task contributes one data point of averaged length on trials that lead to optimal plans (i.e., average-on-optimal length (Sec. 4.1.2)).
(b) Distribution of average-on-optimal length.
(c) Distribution of lengths of individual sequences that lead to an optimal plan.
Figure 8: Improvement of search dynamics length on optimal plans via bootstrapping in Sokoban. After three improvement steps, our Searchformer model reduced the generated search dynamics token length to an average of 3094—a 26.8% reduction in total from 4222, the average $A^{*}$ search length. With each improvement step, the Searchformer models generated shorter and shorter token sequences when computing an optimal plan. 8(a): Comparison of search dynamics length (in terms of number of tokens) between $A^{*}$ search and our models (search-augmented in blue, Searchformer step 1-3 in orange), on the test subset in which our models yield optimal plans. Here, for each test task, we average the lengths of sequences that ends in an optimal plan, out of 64 trials (i.e., average-on-optimal length (Sec. 4.1.2)). The box plots show the distribution of average-on-optimal lengths, in which the left boundary, mid solid line, right boundary represents the 25%, 50% and 75% quantiles. Dotted lines are means and whiskers show the range. 8(b): Histograms of average-on-optimal lengths. 8(c): Histograms of the length of individual sequences that lead to optimal plan. Because it is more likely for each model to generate an optimal plan for easy tasks with fewer search steps, the distributions are skewed towards the left—towards shorter sequence lengths. The data plotted in this figure aggregates sequence lengths across three experiment repeats.
Figure 8 illustrates how the sequence lengths generated by the Searchformer model’s are iteratively shortened by our bootstrapping method. With every improvement step, the length of the generated traces—the number of search steps—decreases (Figure 8(a)). When computing an optimal plan, the final Searchformer model generates search dynamics sequences that are on average 26.8% shorter than the sequences generated with $A^{*}$ search. Consequently, the Searchformer model found a more efficient way to plan a complicated task and go beyond $A^{*}$’s search, even for unseen test puzzles. In Figure 8(b) we can observe that the search-augmented model generates sequences that on average match the sequences generated with $A^{*}$ search in length. The Searchformer models generate shorter sequences resulting in a distribution that is skewed towards shorter sequence lengths.
We find that on average each model produces an optimal plan more often for tasks that can be solved with fewer search steps. Generating an optimal plan for more complex tasks is possible with each model but requires the model to be prompted more often. This is illustrated in Figure 8(c) which plots the sequence length distribution for individual sequences. The sequence length distribution of both models is skewed towards shorter sequences and the models do not match the $A^{*}$ sequence length distribution.
Table 1: Test set performance in the Sokoban tasks. Over 200 unseen test Sokoban tasks, we report percentage of solved, percentage of optimal solutions, Success weighted by cost (SWC) and Improved Length Ratio (ILR) w.r.t $A^{*}$ reference, averaged over sequences ending in an optimal/correct plan. Please check Sec. 4.1 for more detailed definitions of optimality metrics (Sec. 4.1.1) and search length metrics (Sec. 4.1.2). The result shows that continue fine-tuning Searchformer yields shorter search dynamics length and at the same time improved percentage of solved/optimally solved task. All values are averages with errors indicating SEM.
Params. Model Solved (%) Optimal (%) SWC ILR-on-solved ILR-on-optimal
45M Solution only 90.3 $\pm 1.0$ 86.8 $\pm 0.3$ 0.890 $\pm 0.009$
Search augmented 92.5 $\pm 1.0$ 90.8 $\pm 1.6$ 0.924 $\pm 0.011$ 0.908 $\pm 0.020$ 0.919 $\pm 0.019$
Searchformer, step 1 95.5 $\pm 1.0$ 93.5 $\pm 1.0$ 0.953 $\pm 0.010$ 1.054 $\pm 0.025$ 1.062 $\pm 0.015$
Searchformer, step 2 96.0 $\pm 0.5$ 93.4 $\pm 0.6$ 0.957 $\pm 0.005$ 1.158 $\pm 0.025$ 1.181 $\pm 0.012$
Searchformer, step 3 95.5 $\pm 0.8$ 93.7 $\pm 1.6$ 0.953 $\pm 0.009$ 1.292 $\pm 0.044$ 1.343 $\pm 0.067$
175M Solution only 95.7 $\pm 0.2$ 90.0 $\pm 0.8$ 0.949 $\pm 0.003$
Search augmented 95.2 $\pm 0.9$ 93.2 $\pm 1.0$ 0.949 $\pm 0.010$ 0.925 $\pm 0.010$ 0.933 $\pm 0.011$
757M Solution only 96.5 $\pm 0.1$ 92.2 $\pm 1.2$ 0.958 $\pm 0.002$
As reported in Table 1, fine-tuning the model resulted in a significant performance improvement, reducing the rate of incorrect and non-optimal solutions by 40% and 30% respectively. After three steps of fine-tuning, the 45M Searchformer shows similar to a 757M model that learns to predict a plan (task solution) only. The comparison using the more fine-grained SWC measures, which also take sub-optimal plans into consideration, shows a similar trend. See Sec. 4.1 for definitions of evaluation measures.
Reduced length of search dynamics with Searchformer
Lastly, we evaluate to which extend the search-augmented and Searchformer models reduced the number of generated search steps to compute a correct plan. This analysis is different from the results presented in Figure 8 because here we focus on computing length ratios between traces generated with each model and traces generated with $A^{*}$ search using Improved Length Ratio (ILR) measure.
This improvement can be observed in Table 1 for each Searchformer model where with every improvement step the model generates shorter and shorter execution traces leading to higher ILR scores. Furthermore, the scores for the two search-augmented models are slightly below one, indicating that the search-augmented models tend to perform slightly more search steps than $A^{*}$ search, while fine-tuned models has a ILR score higher than 1, showing more efficient search dynamics, in particular for those that lead to optimal plans. For example, $A^{*}$ search dynamics is $\sim 34.3\%$ longer than that produced by Searchformer after 3 steps of fine-tuning. More details about model’s performance characteristics are presented in Appendix A.4.
5 Conclusion, limitations, and future work
Prior work (Momennejad et al., 2023; Valmeekam et al., 2023a) has found that LLMs, which are based on Transformer models, struggle with solving complex decision making tasks. Searchformer demonstrates that with appropriate training data, Transformers can in fact solve complex planning tasks. Moreover, Searchformer robustly follows the intermediate steps—the execution trace—of a symbolic planner and improves (in terms of trace length) beyond the human-crafted rule-based planning strategy it was initially trained on. Compared to solution-only models that directly predict a solution, our search-augmented models require fewer training sequences and scale better to more difficulty tasks.
Limitations
Currently, Searchformer is trained on the execution traces of $A^{*}$ to learn a complex planning strategy. However, the trace length may grow exponentially in the length of an optimal plan (see Figure 3), and training on the resulting token sequence data become computationally very expensive. In fact, the presented experiments use token sequences that are significantly longer than the sequences used to train LLMs such as Llama 2 (Touvron et al., 2023). [^3] The 175 million parameter model trained on the Sokoban dataset uses 64 GPUs with 32GB of memory.
Future work
One way to mitigate this limitation is to use curriculum learning: starting from simple tasks with reasonably long execution traces, train and fine-tune the Searchformer to reduce the trace length, and then adapt the improved model to more complex tasks. Another possibility is to integrate better heuristics or value functions into $A^{*}$ search, similar to Monte Carlo Tree Search (MCTS) (Coulom, 2006), to cap the maximal depth the search algorithm explores. Integrating hierarchical planning methods and temporal abstractions (Sutton et al., 2023, 1999; Dietterich, 2000; Hafner et al., 2022) are another avenue. This would equip the resulting model with the ability to abstract over multiple time steps and states to find an optimal plan using fewer computation steps.
We hope that our study informs further research about how to build Transformer-based models that can robustly solve decision making tasks and mitigate the limitations previously outlined by Momennejad et al. and Valmeekam et al..
6 Broader Impact
Our work focuses on symbolic planning tasks and uses synthetic datasets for training. While the tasks we explored in this paper can be easily solved with simple symbolic solvers, it is important to study the effectiveness of neural networks on such tasks. Neural network based Large Language Models have shown impressive performance on non-symbolic tasks, such as following complex natural language instructions, but still fail on relatively simple symbolic tasks like adding numbers. While the results in this paper are encouraging, they are still far from being competitive with symbolic solvers.
Acknowledgment
We would like to thank Qinqing Zheng for helpful discussions and comments on earlier versions of this paper. We would also like to thank Amy Zhang for helpful discussions on this work.
Appendix A Appendix
A.1 Training encoder-decoder Transformers to predict optimal plans
For this work, we consider encoder-decoder Transformers (Raffel et al., 2020) and process each prompt—the task specification—with an encoder network to obtain an embedding of a planning task. Using the decoder, the resulting embedding is then decoded into a response of the format <trace><plan> or the format <plan>. We refer to a model predicting sequences of the format <trace><plan> as a search-augmented model and a model predicting sequences of the format <plan> as solution-only models.
A.1.1 Encoder-decoder architecture
A token sequence is processed by a neural network by first indexing the set of all tokens contained in a specific dataset. This indexing is used to map any token sequence to a set of integers. Formally, we denote a prompt as a sequence of integers $x_{1:n}=(x_{1},...,x_{n})$. An encoder Transformer neural network with network weights $\boldsymbol{\theta}_{\text{enc}}$ is a function $f_{\boldsymbol{\theta}_{\text{enc}}}$ mapping a tokenized prompt $x_{1:n}$ of arbitrary length $n$ to a sequence of $n$ real-valued vectors:
$f_{\boldsymbol{\theta}_{\text{enc}}}:x_{1:n}\mapsto\boldsymbol{z}_{1:n}$ (1)
Each vector $\boldsymbol{z}_{i}$ in this sequence of vectors $\boldsymbol{z}_{1:n}$ has the same dimension. The decoder network is then used to generate a response auto-regressively: Starting with a specific beginning-of-sequence token bos to cue the decoder, a sequence is recursively built by first mapping the start token bos to a probability distribution over next-tokens. This probability distribution is stored as a vector $\boldsymbol{p}_{1}$ whose dimension is equal to the size of the vocabulary. The next token is then generated by sampling from this distribution and the sampled token is appended to the response sequence. Subsequently the two-token response sequence is fed into the decoder again to compute the next probability vector $\boldsymbol{p}_{2}$ and sample the next token. This procedure is repeated until an end-of-sequence token eos is sampled. While only the last computed probability vector is needed to sample the next token, the decoder network simultaneously predicts a sequence of next token probability vectors $\boldsymbol{p}_{1:m}$ given an input sequence $y_{1:m}$. Furthermore, this prediction is conditioned on the encoder output $\boldsymbol{z}_{1:n}$. The decoder Transformer neural network with weight parameters $\boldsymbol{\theta}_{\text{dec}}$ is therefore a function
$g_{\boldsymbol{\theta}_{\text{dec}}}:\boldsymbol{z}_{1:n},y_{1:m}\mapsto\boldsymbol{p}_{1:m}.$ (2)
The encoder network $f_{\boldsymbol{\theta}_{\text{enc}}}$ and decoder network $g_{\boldsymbol{\theta}_{\text{dec}}}$ internally both use a number of stacked causal attention layers to form an encoder-decoder Transformer (Vaswani et al., 2017) as outlined in more detail in Appendix A.2. We denote a concatenation of all encoder parameters $\boldsymbol{\theta}_{\text{enc}}$ and decoder parameters $\boldsymbol{\theta}_{\text{dec}}$ with the vector $\boldsymbol{\theta}$.
A.1.2 Training with teacher forcing
An encoder-decoder architecture is optimized to generate responses that follow the distribution of a training dataset by minimizing the cross-entropy between the training data distribution $p_{\mathcal{D}}$ over prompt-response pairs $(x_{n},y_{m})$ and the distribution $p_{\boldsymbol{\theta}}$ with which the encoder-decoder model is generating responses. This cross-entropy loss objective
$H(p_{\mathcal{D}},p_{\boldsymbol{\theta}})$ $=\mathbb{E}_{\mathcal{D}}\left[-\log p_{\boldsymbol{\theta}}(y_{1:m}|x_{1:n})\right]$ (3)
$=\mathbb{E}_{\mathcal{D}}\left[-\sum_{i=1}^{m-1}\log p_{\boldsymbol{\theta}}(y_{i+1:m}|y_{1:i},x_{1:n})\right],$ (4)
where line (4) follows from the auto-regressive generation procedure described before. Within the same planning dataset, different prompt-response pairs can have different prompt and response lengths. To emphasize shorter response sequences during training, we re-weigh each sample resulting in the loss objective
$L(\boldsymbol{\theta})=\frac{1}{D}\sum_{d=1}^{D}\frac{1}{m_{d}-1}\sum_{i=1}^{m_{d}-1}\log p_{\boldsymbol{\theta}}(y^{d}_{i+1:m_{d}}|y^{d}_{1:i},x^{d}_{1:n_{d}}),$ (5)
where the first summation averages over all $D$ prompt-response pairs of the training dataset. In Equation (5) the super-script $d$ indexes individual prompt-response pairs in the training dataset. This average is an empirical approximation of the expectation in Equation (3) for a finite i.i.d. sample of size $D$. This loss objective is optimized using gradient descent (Goodfellow et al., 2016, Chapter 10).
A.2 Network architecture and hyper-parameters
Figure 9: Attention blocks architecture used in the encoder-decoder Transformer architecture. The encoder network consists of a set of feed-forward layers where each layer processes hidden activations in parallel with a set of encoder attention blocks (left diagram). Similarly, the decoder network consists of a set of feed-forward layers composed of a number of decoder attention blocks (right diagram). The number of blocks in each layer is referred to as the number of heads. Token sequences are first mapped into integer sequences using a look-up dictionary. Then, these sequences are fed through a PyTorch (Paszke et al., 2019) torch.nn.Embedding layer to map the integer sequence into a sequence of hidden activation vectors. After the last decoder layer, hidden activations are mapped through a linear layer into a sequences of logits vectors to compute the next token probability vectors $\boldsymbol{p}_{1:m}$.
The encoder-decoder Transformer first maps every token to a one-hot vector of the same dimension as the token vocabulary space. These one-hot vectors are then projected through a linear embedding layer into a set of vectors.
The encoder then consists of multiple feed-forward layers and each layer consists of multiple encoder blocks (left part of Figure 9). The output of these layers is then mapped through another linear layer to project the hidden activations into a tensor of the correct shape. The decoder also consists of multiple feed-forward layers and each layer also consists of multiple decoder blocks (right part of Figure 9) processing the hidden activations in parallel. As illustrated in Figure 9, the decoder network is conditioned on the encoder by processing the encoder’s output tensor directly in the second attention map. Furthermore, each tokens position is encoded by applying RoPE embeddings (Su et al., 2023) as indicated in Figure 9. We did not use dropout in our architecture.
Table 2: Architecture Hyper-parameters. The same parameters were used in both the encoder and decoder network. The number of heads indicates how many attention blocks are used in one layer. Layer dimension indicates the dimension of the feature vectors processed through each attention block (dimension of $K$, $V$, and $Q$ in Figure 9). All models used a RoPE frequency of 10000.
Parameter 15M param. model 46M param. model 175M param. model 747M param. model
Layers 6 8 9 16
Heads 3 4 4 12
Layer dim. 64 96 192 96
Table 3: Optimization hyper-parameters. Every model was optimized using AdamW (Loshchilov and Hutter, 2019) with setting $\beta_{0}=0.9$ and $\beta_{1}=0.99$. Initially, the learning rate was linearly interpolated: Starting at zero and then increasing linearly to the value listed below until step 2000. Then a cosine learning rate schedule was followed (Loshchilov and Hutter, 2016). to the between zero at the first training step and the listed value below for the
Parameter Model Maze Tasks Sokoban Puzzels
Learning rate 15M $2.5\cdot 10^{-4}$ $2.5\cdot 10^{-4}$
46M $7.5\cdot 10^{-5}$ $7.5\cdot 10^{-5}$
175M $5.0\cdot 10^{-5}$ $5.0\cdot 10^{-5}$
747M $5.0\cdot 10^{-5}$
Batch size all 16 64
Training steps all 400000 80000
Table 2 lists the used architecture hyper-parameter and Table 3 lists the hyper-parameters used for optimizing each model. All experiments were implemented in PyTorch 2.0 (Paszke et al., 2019) and default parameters were used unless reported here otherwise.
A.3 Dataset generation
bosworker 2 3box 2 4box 3 4dock 1 3dock 4 4wall 0 0wall 0 1...wall 6 6eosPromptboscreate worker 2 3 box 2 4 box 3 4 c0 c3close worker 2 3 box 2 4 box 3 4 c0 c3...create worker 5 4 box 2 3 c10 c3close worker 2 1 c12 c0Trace(2583 tokens)plan 2 3plan 1 3plan 1 4plan 1 5plan 2 5plan 2 4plan 3 4plan 2 4plan 2 3plan 2 2plan 1 2plan 1 1plan 2 1eosPlan
Figure 10: Token sequence example for Sokoban This figure lists the token sequence for the Sokoban level depicted in Figure 2.
(a) Plan Sequence Length
(b) $A^{*}$ Execution Trace Length
Figure 11: Sequence length distribution for each dataset. The training and test sets are designed such that their sequence lengths match closely.
All datasets were generated by first randomly sampling a task and then executing $A^{*}$ to obtain an optimal plan. Maze tasks were generated first by randomly selecting 30-50% of all cells to be wall cells. Then a start and goal location was randomly selected and $A^{*}$ was executed to obtain an optimal plan. If the plan had a length of at least the mazes width or height (e.g. for $10\times 10$ mazes the optimal plan needs to contain at least 10 steps), then the task was added into the dataset. For Sokoban, a $7\times 7$ grid was sampled and two additional wall cells were added as obstacles to the interior of the map. Then two docks, boxes, and the worker locations were randomly placed. If the sampled task is solvable by A*, then the task was admitted to the dataset.
Due to the large volume of generated data, all data was stored in and transformed with a MongoDB (MongoDB Inc., ) instance and map-reduce techniques. Furthermore, when sampling tasks, the dataset was constructed to reject duplicate tasks ensuring that training and test data and each prompt is distinct. Once each task and trace dataset was generated, each execution trace is converted into prompt and response token sequences, as illustrated in Figure 1(c) and Figure 11. Because the Sokoban tasks are very complex and difficult to solve for A*, the resulting token sequences were partially very long and reached almost 100000 tokens. Due to GPU memory requirements, the Sokoban dataset was further sliced to only include sequences of with at most 10000 tokens. Figure 11 compares the sequence length distribution for each dataset. During training, each dataset was also sorted and sliced to only contains the reported number of training sequences. Furthermore, each model was evaluated on the same test tasks within each task dataset. The test dataset contains plans and traces that are of comparable length to the training data (Figure 11).
A.4 Searchformer performance analysis
(a) Comparison of all sequences ending with an optimal plan
(b) Comparison of all sequences ending with a correct plan
(c) Comparison of all sequences ending with a correct but sub-optimal plan
Figure 12: Sequence length comparison between sequences generated with $A^{*}$ search and sequences generated with each model. Each dot in each scatter plot corresponds to one specific test task. On the $x$-axis, we plot the average token sequence length when $A^{*}$ search is used. On the $y$-axis, we plot the average token sequence length when each model is used. Error bars indicate standard deviations. Percentages indicate the fraction of the generated sequences that are included in each scatter plot. 12(a): Sequence length comparison for all test prompts for which an optimal plan was generated using the corresponding model. 12(b): Sequence length comparison for all test prompts for which a correct plan was generated using the corresponding model. This plot aggregates across sequences ending in an optimal plan and sequences ending in a correct but sub-optimal plan. 12(c): Sequence length comparison for all test prompts for which a correct but sub-optimal plan was generated using the corresponding model. This plot only aggregates across sequences ending in a correct but sub-optimal plan. Sequences ending in an optimal plan were withheld.
In Sec. 3.3, each search-augmented and Searchformer model is evaluated by generating 64 token sequences for each Sokoban test task. For the same test task the same model can generate sequences that end in an optimal plan, an incorrect plan, or a correct but sub-optimal plan. In Figure 12 we compare the length of the generated sequences with the length of sequences generated when using $A^{*}$ search for each case. The percentages in each panel’s caption list how many out of all generated sequences end in an optimal plan, a correct plan (that can be optimal or sub-optimal), and a correct but sub-optimal plan.
In the left panel of Figure 12(a) and Figure 12(b), points are centered around the diagonal axis indicating that the search-augmented models do approximately match the $A^{*}$ search algorithm in terms of token sequence lengths. Figure 12(a) and Figure 12(b) further confirm the results presented in Sec. 3.3: With every improvement step, the points move down and below the diagonal line. This highlights that the improved Searchformer models generate token sequences that are shorter than sequences generated with $A^{*}$ search. The Searchformer has found a method of searching a task to compute an optimal plan in fewer search steps than $A^{*}$ search uses.
Figure 12(c) illustrates what happens when each model generates a correct but sub-optimal plan. Here, the search-augmented model, that is trained to imitate $A^{*}$, generates trace sequences that are significantly longer than the sequences generated with $A^{*}$. This suggests that the model struggles in computing a correct plan and generates too many search steps, ultimately leading in finding a correct but sub-optimal plan. Similarly, the Searchformer models also generate sequences that are longer than the sequences generated with $A^{*}$ search. Despite these inefficiencies, our bootstrapping method is still able to improve the model and bring the average sequence length closer to the length of sequences generated with $A^{*}$ search (right most panel in Figure 12(c)). While we would desire the trace length to be low in either case, we found that each model generates a correct but sub-optimal plan with less than 5% chance. Consequently, Figure 12(b) shows that the final Searchformer model still generates a correct plan with on average fewer search steps than $A^{*}$ search. Statistically, the differences between Figure 12(a) and Figure 12(b) are marginal.
A.5 Supplemental maze experiments
(a) Deterministic Plan Prediction
(b) Non-deterministic Plan Prediction
Figure 13: Optimal plan prediction performance for solution-only models. Each panel plots the percentage of correctly answered prompts averaged across five different seeds.
(a) Deterministic Prediction
(b) Non-deterministic Prediction
Figure 14: Optimal plan prediction performance for search-augmented models. Training an encoder-decoder Transformer to imitate A* execution traces significantly boosts performance. 14(a): Percentage of correctly generated responses. Note that the plan only models only need to generate a few hundred token long plan to answer a prompt correctly. The trace plan model must generate a much longer A* execution trace correctly to correctly answer a prompt. This requires the trace plan models to generate response sequences that are hundreds or thousands of tokens long (cf. Figure 3). If one token is incorrectly predicted, the response is counted as incorrect. 14(b): Percentage of correctly generated plans. Each model was used to generate multiple responses for each prompt. If one of the generated responses contains an optimal plan, the test prompt is counted as correctly answered.
Figure 13 and Figure 14 plot how often a correct answer was provided by each model for all studied maze shape, training dataset size, and model parameterization combinations.

### v2

@@
 Abstract
+While Transformers have enabled tremendous progress in various application settings, such architectures still trail behind traditional symbolic planners for solving complex decision making tasks.
+In this work, we demonstrate how to train Transformers to solve complex planning tasks.
+This is accomplished by training an encoder-decoder Transformer model to predict the search dynamics of the $A^{*}$ search algorithm.
+We fine tune this model to obtain a Searchformer, a Transformer model that optimally solves previously unseen Sokoban puzzles 93.7% of the time, while using up to 26.8% fewer search steps than the $A^{*}$ implementation that was used for training initially.
+In our training method, $A^{*}$’s search dynamics are expressed as a token sequence outlining when task states are added and removed into the search tree during symbolic planning.
+Searchformer significantly outperforms baselines that predict the optimal plan directly with a 5–10$\times$ smaller model size and a 10$\times$ smaller training dataset.
+Lastly, we demonstrate how Searchformer scales to larger and more complex decision making tasks with improved percentage of solved tasks and shortened search dynamics.
 […]
-[…]
+[Code]https://github.com/facebookresearch/searchformer
 1 Introduction
-Over the past few years, Transformer-based architectures (Vaswani et al., 2017) have demonstrated impressive performance in different tasks, including holding conversations at the human level (Shuster et al., 2022; OpenAI, 2022, 2023; Touvron et al., 2023), high-quality image understanding (Caron et al., 2021; Oquab et al., 2024; Assran et al., 2023) and video generation (Singer et al., 2023), multi-modal generation (Girdhar et al., 2023; Radford et al., 2021), and code completion (Roziere et al., 2023; OpenAI, 2021).
-By training these architectures with a huge amount of data, the resulting models, such as Large Language Models (LLMs), can generalize well in real-world use cases.
+Transformer-based architectures (Vaswani et al., 2017) have demonstrated impressive performance in different tasks, including holding conversations at the human level (Shuster et al., 2022; OpenAI, 2022, 2023; Touvron et al., 2023), high-quality image understanding (Caron et al., 2021; Oquab et al., 2024; Assran et al., 2023) and video generation (Singer et al., 2023), multi-modal generation (Girdhar et al., 2023; Radford et al., 2021), and code completion (Roziere et al., 2023; OpenAI, 2021).
+By training these architectures on internet-scale datasets, the resulting models, such as Large Language Models (LLMs), can generalize well in real-world use cases.
 Despite these successes, Transformer-based architectures and LLMs still struggle when it comes to solving planning and reasoning tasks.
-Previous studies (Momennejad et al., 2023; Valmeekam et al., 2023a, b) demonstrate that LLMs fall short in multi-step planning tasks (Valmeekam et al., 2023b) or when performing higher-order reasoning (Momennejad et al., 2023; Fan et al., 2020).
-In recent years, various methods have been proposed to improve the performance of Transformers in reasoning and planning tasks.
-One common and effective approach is to simulate the human thinking process and produce intermediate “thoughts” before outputting a response.
-Chain-of-Thought (CoT) prompting (Wei et al., 2022) encourages the model to predict the intermediate steps and to “think” step by step.
-Tree-of-thoughts (ToT) uses a branching strategy and critics to generate different thought paths before picking the best one (Yao et al., 2023).
-While these techniques are often effective, there are studies showing that in many cases, they may lead to worse performance, for example due to self-enforcing (Huang et al., 2023).
+Previous studies demonstrate that LLMs fall short in multi-step planning tasks (Valmeekam et al., 2023a, b) or when performing higher-order reasoning (Momennejad et al., 2023; Fan et al., 2020).
+In recent years, various methods have been proposed to improve the performance of Transformers in these settings.
+One approach is to simulate the human thinking process and produce intermediate “thoughts” before outputting a response.
+Chain-of-Thought (CoT) prompting (Wei et al., 2022) and the Tree-of-thoughts (ToT) method (Yao et al., 2023) encourage the model to “think” step by step.
+While these techniques are often effective, they can also lead to worse performance, for example due to self-enforcing (Huang et al., 2023).
 Furthermore, techniques effective on one dataset may not work well on others due to changes in the type of reasoning involved (e.g., spatial reasoning vs.
-mathematical reasoning vs.
-common-sense reasoning).
-How to enable Transformers or LLMs to plan, solve multi-step decision making tasks, and perform different types of reasoning still remains elusive and an active area of research.
-These methods stand in sharp contrast with traditional symbolic planning and search techniques.
-While such techniques may not exhibit the language understanding capabilities of LLMs trained on internet-scale datasets, they are well-defined and can solve planning tasks that are far more complex than the tasks state-of-the-art LLMs can solve (Momennejad et al., 2023; Valmeekam et al., 2023a).
-Additionally, there often exist formal guarantees for solutions computed by these traditional methods, because symbolic planning algorithms often follow a well-defined rule-based search procedure.
+mathematical reasoning).
+How to enable Transformers and LLMs to plan, solve multi-step decision making tasks, and perform reasoning still remains elusive and an active area of research.
 Our work
-In this paper, we demonstrate how to train Transformers to solve complex planning tasks and present Searchformer, a Transformer model that computes an optimal plan in fewer search steps than symbolic planning algorithms such as $A^{*}$ search (Russell and Norvig, 2021, Chapter 3) in multi-step planning tasks such as maze navigation and solving Sokoban puzzles.
-To achieve this, we introduce search dynamics bootstrapping, a method that first trains a Transformer model to imitate the search procedure of $A^{*}$ and then fine-tunes the model to find an optimal plan with fewer search steps.
-In the first step a Transformer model is trained to imitate $A^{*}$ search.
-Here, we generate synthetic datasets by running $A^{*}$ search in randomly generated planning task instances.
-While $A^{*}$ is executed, we log the performed computation and optimal plan into a sequence of words, called tokens.
-The resulting training dataset then contains an execution trace of $A^{*}$ and encodes information about the search dynamics of $A^{*}$ itself.
-A Transformer model is then trained to generate these token sequences along with an optimal plan for any given planning task.
-In the second step, we show that our Searchformer trained with search-augmented sequences—sequences that include $A^{*}$’s execution trace—can be further improved via expert iteration (Gulcehre et al., 2023), a method that encourages the Transformer to generate fewer search steps while still generating optimal solutions.
-This procedure leads to a neural planning algorithm that is implicitly encoded in the Transformer’s network weights, and finds an optimal plan with high probability in fewer search steps than standard $A^{*}$ search.
-Specifically, when solving Sokoban puzzles, our models solve 93.7% of all test tasks while performing search steps that are on average 26.8% shorter than $A^{*}$ search.
-This paves the way to use Transformers to move beyond traditional symbolic planning algorithms.
-To better understand how the used training data and number of model parameters influence performance of the resulting models, we present a set of ablation studies.
-We train models on both solution-only sequences (token sequences including task description and final plan) and search-augmented sequences (token sequences including task description, search tree dynamics, and final plan) and compare their performance.
-Despite the fact that search-augmented sequences are thousands of tokens long, search-augmented models (that are trained on search-augmented sequences) can still learn to generate the correct search dynamics and optimal plans with high probability.
-Further, compared to solution-only models (that are trained on solution-only sequences), search-augmented models generate an optimal plan more often on unseen tasks with ten times fewer training sequences, highlighting the power of including $A^{*}$’s search dynamics into the training process of Transformer models.
-Since the search dynamics can be included into the training data simply by through logging when generating synthetic training sequences, our study provides a data-efficient way to improve the reasoning power of Transformer models.
-We also find that solution-only models do not improve with more parameters, showing that the type of training data used plays a critical role in this setting.
+We demonstrate how to train Transformers to robustly solve complex planning tasks.
+Similar to LLMs, we train Transformers to predict the next word given a sequence of words.
+Our experiments use synthetically generated datasets with a synthetic language and vocabulary.
+Using this framework, we demonstrate how to construct training data such that the resulting model imitates the computation performed by $A^{*}$ search (Russell & Norvig, 2021, Chapter 3).
+Lastly, we present Searchformer, a Transformer model that solves complex planning tasks in fewer search steps than our $A^{*}$ reference implementation.
+This model is obtained through search dynamics bootstrapping, a method that first trains a Transformer to imitate $A^{*}$’s search process and then fine-tunes the model to find an optimal plan within fewer search steps.
+To train a Transformer to perform planning, we express a planning task and its optimal solution plan as a sequence of words, called tokens.
+We also log the computation performed by $A^{*}$ into an execution trace consisting of a token sequence, resulting in a sequence dataset that captures $A^{*}$’s search dynamics.
+Using these search-augmented sequences, a Transformer model is trained to generate token sequences that encode $A^{*}$’s search dynamics along with an optimal plan.
+Subsequently, the resulting search-augmented model is fine-tuned to generate shorter token sequences while still outputting an optimal plan.
+We refer to this final fine-tuned model as a Searchformer.
+When solving Sokoban puzzles, our models solve 93.7% of all test tasks while performing on average 26.8% fewer search steps than our $A^{*}$ reference implementation.
+Through a sequence of experiments that control task complexity, dataset size, and model size, we demonstrate that including execution traces into the training data increases performance on an independent test task set—despite a 10×10$\times$10 × to 100×100$\times$100 × increase in the length of the generated sequences.
+We find that search-augmented models (that include execution traces into their training data) generate an optimal plan more often on unseen tasks with ten times fewer training sequences than a larger solution-only model (that is trained on sequences only including a task description and task solution).
+This result highlights the power of including $A^{*}$’s search dynamics into the training process of Transformer models.
 2 Related Work
-While existing work (Trinh et al., 2024; Ruoss et al., 2024) leverages synthetic datasets to learn strong policies for reasoning, our study is fundamentally different in this regard.
-We focus on improving the reasoning capability embedded in a Transformer’s weights, while algorithms such as AlphaZero (Silver et al., 2018), MuZero (Schrittwieser et al., 2020) and AlphaGeometry (Trinh et al., 2024) treat general neural network models (including Transformers) as black-boxes, and use existing symbolic planning techniques to improve the neural network itself.
+While existing work (Trinh et al., 2024; Ruoss et al., 2024) leverages synthetic datasets to learn policies for reasoning, our study is fundamentally different in this regard.
+We focus on improving the reasoning capability embedded in a Transformer’s weights.
+Existing algorithms such as AlphaZero (Silver et al., 2018), MuZero (Schrittwieser et al., 2020), and AlphaGeometry (Trinh et al., 2024) optimize a neural network using the output of existing symbolic planning algorithms, whose internal states are not used (i.e., treated as black-boxes).
 For example, Silver et al.
 (2017) use MCTS as a policy improvement operator to update the neural network’s weights.
-In contrast, the search dynamics bootstrapping method in this work uses a Transformer model to generalize towards more efficient search patterns and improve the model itself.
-A planning algorithm is only used to initially train a Transformer model.
-ScratchPad (Nye et al., 2021) train neural networks on execution traces of reasoning tasks, Yang et al.
-(2022), Pallagani et al.
-(2022), and Ruoss et al.
-(2024) train models to predict the optimal solution or an optimal action directly.
-In contrast, we find more efficient LLM-based planning algorithms by predicting the search dynamics of a planning algorithm, which is the execution trace when the solver attempts to find the optimal solution, which is a harder problem than predicting an optimal decision making strategy.
-Our work also bears some similarity with neuro-symbolic systems (e.g., Neural Turing Machines (Graves et al., 2014), Neural Programming Architecture (Cai et al., 2017)), which build differentiable architectures to mimic the functionality of existing symbolic systems.
-However, these methods use dedicated components (e.g., explicit memory components, built-in recursion), while Searchformer uses a general Transformer architecture.
-By using long contexts and positional embeddings (Chen et al., 2023; Peng et al., 2023) to predict an optimal plan, Searchformer performs well in solving different complex planning tasks.
-Ultimately, our work is a step towards building more general neural architectures that automatically learn a reasoning and planning mechanism without prior human knowledge.
+In contrast, the presented search dynamics bootstrapping method uses a Transformer model to generalize towards more efficient search patterns and improves the model itself.
+A planning algorithm (together with its internal search dynamics) is used to initially train a Transformer model.
+Prior work focuses on training a neural network on execution traces of reasoning tasks (Nye et al., 2021) or on training a neural network to predict an optimal action (Yang et al., 2022; Pallagani et al., 2022; Ruoss et al., 2024).
+In contrast, we focus on training a Transformer to generate the entire search process of $A^{*}$ when computing an optimal plan.
+Instead of only predicting a single action, our model predicts an entire multi-step plan to solve a task.
+Our work bears some similarity with neuro-symbolic systems (Graves et al., 2014; Cai et al., 2017), which build differentiable architectures to mimic the functionality of existing symbolic systems.
+However, these methods use dedicated components (e.g., explicit memory components, built-in recursion), while Searchformer focuses on next-token prediction.
+Here, Searchformer relies on generating long contexts and position embeddings (Chen et al., 2023; Peng et al., 2023) to predict in optimal plan.
+Ultimately, our work sheds light on how to build more general architectures that automatically learn a planning mechanism.
+Using Transformer architectures to solve complex sequential decision making tasks has been studied in prior work in a reinforcement learning (RL) setting (Chen et al., 2021; Janner et al., 2021; Laskin et al., 2023).
+However, this prior work presents different methods to modelling trajectories of trial and error interactions and focuses on predicting a next action, state, or rewards or a combination of them.
+In contrast, we demonstrate how to use a Transformer to model the search steps involved in computing an optimal multi-step plan.
+MCTSNet (Guez et al., 2018) also attempts to learn the search procedure itself, but still hard-codes the MCTS search procedure (Coulom, 2006) into a neural network, which leads to quadratic backpropagation overhead and can only deal with up to 500 step rollouts, while our approach can deal with much longer search execution trace.
+We demonstrate that Transformers can not only imitate a symbolic planning algorithm but can also be used to discover more efficient heuristics via fine-tuning.
 : wall cell: start cell: goal cell: plan step210210
 (a) Maze navigation task
-Promptbosstart 0 2goal 1 0wall 1 2wall 2 0eosResponsebosplan 0 2plan 0 1plan 1 1plan 1 0eos
+Promptbosstart  0 2goal   1 0wall   1 2wall   2 0eosResponsebosplan  0 2plan  0 1plan  0 0plan  1 0eos
 (b) Tokenization of a planning task and its solution
-[…]
+A* planning algorithm 0: Start node $n_{\text{start}}$ and goal node $n_{\text{goal}}$.
+1:$\mathcal{S}_{\text{closed}}\leftarrow\{\}$ 2:$\mathcal{S}_{\text{frontier}}\leftarrow\{n_{\text{start}}\}$ 3:while  $|\mathcal{S}_{\text{frontier}}|>0$  do 4:    $n_{\text{curr}}=\arg\min_{n\in\mathcal{S}_{\text{frontier}}}\text{cost}(n)$ 5:    $\mathcal{S}_{\text{closed}}\leftarrow\mathcal{S}_{\text{closed}}\cup\{n_{\text{curr}}\}$ 6:    for  $n_{\text{child}}\in\text{children}(n_{\text{curr}})$  do 7:       if $\text{pos}(n)=\text{pos}(n_{\text{child}})$ for any $n\in\mathcal{S}_{\text{closed}}\cup\mathcal{S}_{\text{frontier}}$ then 8:          if $c(n)+h(n)\leq c(n_{\text{child}})+h(n_{\text{child}})$ then 9:             continue 10:          end if 11:       end if 12:       Set $\text{parent}(n_{\text{child}})\leftarrow n_{\text{curr}}$ 13:       $\mathcal{S}_{\text{frontier}}\leftarrow\mathcal{S}_{\text{frontier}}\cup\{n_{\text{child}}\}$ 14:    end for 15:end while 16: Compute and return plan by recursing on parents of $n_{\text{curr}}$.
+Tokenization of algorithm executionAdd node to frontierAdd node to closed setCost from start $c(n)$Heuristic value $h(n)$boscreate 0 2 c0 c3close  0 2 c0 c3create 0 1 c1 c2close  0 1 c1 c2create 0 0 c2 c1create 1 1 c2 c1close  0 0 c2 c1create 1 0 c3 c0close  1 0 c3 c0plan    0 2plan   0 1plan   0 0plan   1 0eosTracePlan
 (c) $A^{*}$’s execution when solving a planning task is logged into an execution trace
 Figure 1: Expressing a planning task in token sequences.
-1(a): A $3\times 3$ maze navigation task where an agent must navigate from the start cell to the goal cell without entering a wall cell.
-For each transition, a cost of one is given and the optimal plan is the shortest path from start to goal.
+1(a): A $3\times 3$ maze navigation task where the goal is to find a the shortest path from start to goal without entering a wall cell.
 1(b): The $3\times 3$ maze navigation task is expressed as a prompt token sequence (left panel) and the optimal plan is expressed as a response token sequence (right panel).
-The start and end of a response sequence is indicated by a beginning-of-sequence token, bos, and an end-of-sequence tokens, eos.
-The numbers indicate $x,y$ coordinates.
-1(c): The search dynamics of the $A^{*}$ algorithm (left panel) is logged into an execution trace represented by a token sequence (right panel).
-For maze tasks, coordinates are encoded with integer tokens.
-The next two numbers costs associated with that node (letter “c” is added to differentiate them from coordinate numbers).
-Each token is mapped to a unique non-negative integer and this integer is processed in the neural network.
-Using Transformer architectures to solve complex sequential decision making has been studied in prior work in a reinforcement learning (RL) setting (Chen et al., 2021; Janner et al., 2021; Laskin et al., 2023).
-Decision Transformer (Chen et al., 2021) predicts the next optimal action given a trajectory segment that is annotated with total return values, while Trajectory Transformer (Janner et al., 2021) focuses on learning a world model that predicts future states, actions, and rewards.
-Similarly, in Algorithm Distillation (Laskin et al., 2023) a Transformer architecture is trained to predict the next action given a trajectory segment of trial and error interactions as context.
-In contrast, we incorporate how a planning algorithm arrives at a final result into a Transformer models.
-MCTSNet (Guez et al., 2018) also attempts to learn the search procedure itself, but still hard-codes the MCTS search procedure into neural network, which leads to quadratic backpropagation overhead and can only deal witih up to 500 rollouts, while our approach can deal with much longer search execution trace.
-We demonstrate that Transformers can not only imitate a symbolic planning algorithm but also improve past the algorithm it was initially trained on via fine-tuning.
+The start and end of a sequence is indicated by a beginning-of-sequence token, bos, and an end-of-sequence token, eos.
+Numbers indicate $x,y$ coordinates.
+1(c): The search dynamics of the $A^{*}$ algorithm (left panel) is logged into an execution trace (right panel).
+The last two tokens in the trace encode the cost-since-start value $c(n)$ and the heuristic value $h(n)$ (letter “c” distinguishes costs from coordinate numbers).
+The $A^{*}$ algorithm is described in detail by Russell & Norvig (2021, Chapter 3).
 3 Problem Setup
-Figure 2: Example Sokoban puzzle [^2] This level image was composed using image icons from https://github.com/morenod/sokoban (accessed 2023-11-21)..
-The prompt, $A^{*}$ execution trace, and optimal plan for this task is illustrated in Figure 10 in Appendix A.3.
-For Sokoban, we only use a non-deterministic $A^{*}$ implementation.
-We consider two domains: maze navigation tasks (Figures 1(a)) and solving Sokoban puzzles (Figure 2).
-In maze navigation, the goal is to find the shortest path in an $n$-by-$n$ maze between a start location and a goal location without passing through wall cells.
-Figure 2 depicts a Sokoban puzzle where the worker can move up, down, left, or right to push each box onto a dock (crossed grid cells).
-The worker cannot pull a box away from a wall cell.
-To solve the puzzle, each box must be moved onto a dock in as few steps as possible.
-An incorrect move may immediately lead to a dead end and thus careful planning is needed to solve the task.
-Furthermore, each state in a Sokoban puzzle consists of every possible combination of a box and worker position.
-Consequently, state spaces are large making a Sokoban puzzle computationally more difficult to solve for algorithms such as $A^{*}$ search.
+Figure 1 provides an overview of our synthetic dataset generation process.
+We consider two domains: maze navigation (Figure 1(a)) and solving Sokoban puzzles (Figure 5 in Appendix C).
+In maze navigation, the goal is to find the shortest path through an $n$-by-$n$ maze.
+In Sokoban, a worker can move up, down, left, or right and has to push each box onto a dock to solve the puzzle.
+An incorrect move may immediately lead to a dead end and thus reasoning across multiple time steps is required to solve the puzzle.
+Each state in a puzzle consists of a combination of box and worker positions, making Sokoban computationally more difficult to solve than maze navigation.
 3.1 Generating execution traces of $A^{*}$ search.
 The $A^{*}$ algorithm computes an optimal plan by manipulating two sets of nodes:
-  • A frontier set containing the current search frontiers;
+  • A frontier set containing the current search frontiers.
   • A closed set containing all searched nodes.
@@
 At any given iteration, which node is searched next is determined by the content of the frontier and closed sets as well as the heuristic and cost from start values (Figure 1(c), left panel).
-The $A^{*}$’s execution trace is collected by tracking all insertion operations into the frontier and closed set along with heuristic and cost from start values.
+$A^{*}$’s execution trace is collected by tracking all insertion operations into the frontier and closed set along with heuristic and cost from start values (Figure 1(c), right panel).
 The right panel in Figure 1(c) illustrates the resulting trace for the maze example shown in Figure 1(b).
@@
 The resulting plan is then appended to this trace.
-This trace is constructed such that given any prefix of this trace the next token can be predicted correctly.
+This trace is constructed such that given any prefix the next token can be predicted correctly.
 For the maze datasets, $A^{*}$ uses the Manhattan distance to the goal location as a heuristic.
@@
 For each experiment, we generate two token sequence variants, as illustrated in Figure 1:
-  • Search-augmented sequence. This variant contains sequences of the format <prompt><trace><plan>, where the <trace> part encodes $A^{*}$ execution trace and <plan> encodes the optimal plan.
-  • Solution-only sequence. This variant omits $A^{*}$’s exeuction trace and only contains sequences of the format <prompt><plan>.
-For example, the $3\times 3$ maze navigation task (Figure 1(a)) can be mapped into a prompt and plan sequence (Figure 1(b)).
-The <trace> part is constructed by running $A^{*}$ on each planning task and logging the performed computations into a token sequence (Figure 1(c)).
-While the optimal plan in both cases can be fairly short, the corresponding execution traces can be quite long, in particular when $A^{*}$’s heuristic underestimates the cost to reach the goal state and often misleads the search to explore in-feasible paths.
-As illustrated in Figure 1, all token sequences are synthetically generated and use a special vocabulary.
-Furthermore, every model is trained from scratch—no natural language dataset is used for training.
-Consequently, the resulting models are specifically trained to only predict sequences that outline optimal plans for a set of different planning tasks.
-After training, the model’s output is parsed and evaluated if it contains a correct and optimal solution plan.
+  • Solution-only sequences of the format <prompt><plan>, where the <prompt> part encodes a task description and the <plan> part the optimal plan (Figure 1(b)).
+  • Search-augmented sequences of the format <prompt><trace><plan>, where the <trace> part encodes $A^{*}$’s execution trace (Figure 1(c)).
+Because every model is trained from scratch, the resulting models are specifically trained to only predict sequences that outline optimal plans for a set of different planning tasks.
+After training, the model’s output is parsed and evaluated if it contains an optimal or feasible solution plan.
 3.2 Training a Transformer model
-To train a Transformer model, we randomly generate a set of planning tasks together with their corresponding token sequences.
-In each dataset, each task is unique and the test set is constructed such that it does not contain any duplicate of the training set.
-For example, a maze task that occurs in the training dataset does not occur in the test dataset and no task is repeated.
-With this experiment design, we hope to gain insight into how Transformers can be used to solve planning tasks.
+When generating a token sequence dataset, each task is unique and the test set is constructed such that it does not contain any duplicate of the training set.
+With this experiment design, we hope to gain insight into how Transformers can be used to solve planning tasks and generalize to previously unseen test tasks.
 By including intermediate computation steps, the Transformer model is trained to effectively imitate the computation performed by the $A^{*}$ algorithm.
-This approach to training a neural network is similar to Procedure Cloning (Yang et al., 2022) where a neural network is not only presented with input-output pairs (in our case task prompts and optimal plans), but also with examples of how to “think” about a particular task
-In our experiments, we train an adaptation of the encoder-decoder T5 architecture (Raffel et al., 2020) and use Rotary Position Embeddings (RoPE) (Su et al., 2023).
-More details and hyper-parameters about the models we use can be found in Appendix A.2.
-The encoder processes the <prompt> part of a training sequence, and the decoder processes either a <trace><plan> formatted sequence to train a search-augmented model or only a <plan> formatted sequence when training a solution-only model.
+Different from Procedure Cloning (Yang et al., 2022) where a neural network is learned to predict the optimal state/action sequence (in our case task prompts and optimal plans), our Transformer model also learns to predict the entire thinking process, including the attempted but failed paths, that leads to the optimal plan.
+For each experiment an adaptation of the encoder-decoder T5 architecture (Raffel et al., 2020) is used that integrates Rotary Position Embeddings (RoPE) (Su et al., 2023).
+More details and hyper-parameters can be found in Appendix B.
+The encoder processes the <prompt> part of a training sequence, and the decoder processes either a <trace><plan>-formatted sequence (search-augmented model) or only a <plan>-formatted sequence (solution-only model).
 Depending on the model variant, each network is trained to maximize the cross-entropy between the distribution of the decoder generations and the distribution of sampling a corresponding sequence from the training dataset.
-Gradient updates are computed using teacher forcing.
-Appendix A.1 describes our optimization setup in more detail.
+Appendix A describes our optimization setup in more detail.
 3.3 Moving past algorithm imitation via search dynamics bootstrapping
-To reduce the number of search steps generated by a search-augmented model during inference, we implement a method to shift the distribution with which the decoder generates execution traces.
-First, a search-augmented model is trained to imitate the search dynamic of our non-deterministic $A^{*}$ implementation.
-This non-deterministic $A^{*}$ implementation breaks cost ties randomly and randomizes the order with which child nodes are expanded.
-By doing so, we do not decrease the efficiency of $A^{*}$ search itself and merely change the order with which different nodes are searched while still respecting $A^{*}$’s heuristic and cost calculations.
-Using a non-deterministic implementation randomizes the sequence generation process and induces additional variance into the length of each execution trace.
+To reduce the number of tokens generated by a search-augmented model during inference, we implement a method to shift the distribution with which the decoder generates execution traces.
+First, a search-augmented model is trained to imitate the search dynamics of $A^{*}$ search.
+To discover new search dynamics with this search-augmented model and explore the execution trace space, the search-augmented model must generate different sequences for the same task prompt.
+We accomplish this by inducing non-determinism into the training data and use a non-determinsitic $A^{*}$ implementation that breaks cost ties randomly and randomizes the order with which child nodes are expanded.
+This approach does not decrease the efficiency of $A^{*}$ search itself and merely changes the order with which different nodes are searched while still respecting $A^{*}$’s heuristic and cost calculations.
 The resulting search-augmented model will then approximate the probability distribution with which the training sequences were generated.
 Once a model is trained to imitate the search dynamics of non-deterministic $A^{*}$ search, it is used to generate a new training dataset consisting of shorter token sequences.
-This new dataset is constructed by using the trained search-augmented model to generate multiple token sequences for each training prompt.
+This new dataset is constructed by using the trained search-augmented model to sample multiple different token sequences for each training prompt.
 In this step, we only use the training dataset for bootstrapping and not the test dataset.
@@
 If this is the case and the sequence is also shorter than the corresponding sequence contained in the original training dataset, then this shortened sequence is included in the new short sequence training dataset.
-If the generated sequence does not end in an optimal plan or is longer, then the sequence from the original training dataset is re-used.
+If the generated sequence does not end in an optimal plan or is longer than the original training sequence, then the sequence from the original training dataset is re-used.
 Subsequently, the search-augmented model is fine-tuned on the new short sequence training dataset.
-To distinguish from the search-augmented model trained on the original dataset, we call this new model Searchformer.
-This procedure can then be repeated by using the resulting fine-tuned model to generate the next even shorter sequence dataset and then using this dataset to further fine-tune the Searchformer model.
-In Section 4.4 we demonstrate that this procedure does in fact reduce the number of search steps performed during inference while further improving performance.
-At this point, the Searchformer model no longer imitates $A^{*}$ search and has instead discovered a new way of solving a planning problem using fewer search steps.
+To distinguish from the search-augmented model that imitates $A^{*}$’s search dynamics, we call this new model Searchformer.
+This procedure can then be repeated by using the resulting fine-tuned model to generate the next even shorter sequence dataset and then fine-tuning the Searchformer model again.
+In Section 4.3 we demonstrate that this procedure does in fact reduce the number of steps performed during inference while further improving performance.
+The Searchformer model no longer imitates $A^{*}$ search and has instead discovered a new way of solving a planning problem using fewer steps.
 4 Experiments
-In our experiments, we use a deterministic and a non-deterministic variant of $A^{*}$ search to generate each sequence dataset.
-Deterministic $A^{*}$ dataset
-In the first case, sequences are generated by executing $A^{*}$ in a deterministic fashion (by ordering child nodes and breaking equal cost ties in the frontier set deterministically).
+In our experiments, we use two different $A^{*}$ implementations for sequence data generation:
+  1. 1.
+    Deterministic $A^{*}$ dataset: Sequences are generated by executing $A^{*}$ in a deterministic fashion (by ordering child nodes and breaking equal cost ties deterministically).
 Consequently, given a task prompt, the optimal plan and $A^{*}$ execution trace is unique.
-A Transformer model trained on this data only has to learn which token to predict next.
-Evaluating such a model is simple, because the generated token sequences need to exactly match the token sequences generated by $A^{*}$.
-Non-deterministic $A^{*}$ dataset
-Figure 3: Training Sequence Length Comparison.
-The left panel plots the length of the solution-only sequences and the right panel plots the length of the search-augmented sequences, excluding the start and end of sequence tokens bos and eos.
-The whiskers indicate the range of all sequence lengths and the box plots indicate the 25%, 50%, and 75% quantiles.
-Because we focus on planning in complex sequential decision making tasks, the token sequences are multiple orders of magnitude longer than usual token sequences used to train LLMs—especially when $A^{*}$ execution traces are included in the responses.
-For example, fine-tuning of the Llama 2 model on human preference data is performed with sequences that are on average 600 tokens long (Touvron et al., 2023).
-In practice, there can be multiple plans that solve the same task optimally and a deterministic implementation of $A^{*}$ search will only compute one optimal solution.
-To diversify the dataset and include different search sequences, we generate sequences by executing $A^{*}$ search in a non-deterministic fashion: in each run, we randomly shuffle the order with which child nodes are expanded and break equal cost ties randomly.
-Consequently, the generated <trace><plan>-formatted sequences vary between different executions.
-The resulting plans may differ as well but are still optimal.
-This modification does not decrease the efficiency of $A^{*}$ search, but merely forces our implementation to use different search sequences that still correspond to executing the pseudo-code listed in Figure 1(c).
-Other randomization strategies—which do influence $A^{*}$’s efficiency—are also possible (e.g., randomizing cost values at each step) and are left for future works.
-Statistics
-Figure 3 presents an overview of the token sequence length used for each model type for the Sokoban dataset and each maze datasets.
-The length of the generated $A^{*}$ execution traces grows with the grid size of each maze.
-In Sokoban, the task with the largest state space, execution traces are the longest.
-However, the solution plans for Sokoban are amongst the shortest.
-4.1 Evaluation Criteria
-For each task, one model may generate a search sequence ending either in an optimal plan, a correct but sub-optimal plan, or an invalid plan.
-For a plan to be valid, we not only check whether the final solution sequence is feasible, but also verify the output format.
-The search-augmented models must generate a token sequence of the format <trace><plan>, while solution-only models must generate a token sequence of the format <plan> only.
-Generated sequences violating such constraints—including sequences that do not predict an end-of-sequence token eos—are omitted.
-4.1.1 Measuring plan optimality
-We evaluate whether the output plan is optimal using one of three criteria:
-  • Exact-match criterion.
-For each task, if the generated sequence from a trained model matches the output of deterministic $A^{*}$ exactly, it is labelled as correct, otherwise labelled as incorrect.
-This is only used to evaluate supervised cloning of deterministic $A^{*}$.
-  • Any-optimal-64 criterion.
-For each task, we sample 64 responses from a trained model.
-They are parsed and then evaluated to test if the final plan is feasible and optimal, regardless of the generated <trace> part.
-If any of the 64 plans is feasible and optimal, then the task is labelled as correct.
-  • SWC score.
-To further measure the sub-optimalness of the resulting plans, we also report the Success Weighted by Cost (SWC) (Wu et al., 2019), a statistic that factors in how close the cost $l_{i}$ of the best predicted correct plan (over 64 trials) is to the optimal plan cost $l_{i}^{*}$, averaged across all $n$ test tasks and weighted by a binary variable $c_{i}\in\{0,1\}$:
-    $\text{SWC}:=\frac{1}{n}\sum_{i=1}^{n}c_{i}\frac{l_{i}^{*}}{\max\{l_{i},l_{i}^{*}\}}.$
-    When computing the SWC score, the binary variable $c_{i}$ is set to one if a correct plan was found and zero otherwise.
-This SWC score lies between zero and one.
-If all generated sequences end in an optimal plan, then this value is one.
-4.1.2 Measuring search dynamics length
-Given the plan is correct or optimal, we evaluate the length of sequence dynamics in terms of number of tokens using one of the two criteria:
-  • Average-on-optimal length. For each task, we sample 64 responses from a trained model and compute averaged search dynamics length for sequences that lead to an optimal plan.
-  • ILR score.
-To further measure how much improvement of the model-generated search dynamics against $A^{*}$ planning, we report the Improved Length Ratio of Search Dynamics (ILR) score.
-Specifically, for each test task $i$, we compute the ratio between the length $t_{i}$ of the shortest generated search dynamics sequence and the $A^{*}$ token sequence length $t_{i}^{*}$.
-We then average across all test tasks while only including ratios for tasks for which either an optimal or a correct (and potentially sub-optimal) plan was found, as specified by the binary variable $c_{i}$.
-The corresponding measures are thus called ILR-on-optimal and ILR-on-solved.
-The ILR is defined as
-    $\text{ILR}:=\frac{1}{n}\sum_{i=1}^{n}c_{i}\frac{t_{i}^{*}}{t_{i}}.$
-    The ILR measure can take non-negative values and values above one indicate that the model generates shorter search dynamics than the $A^{*}$ reference.
-Consequently, if the numbers lie significantly above one, then the model has found a more efficient way to search a task’s state space to compute an optimal plan.
-Unless indicated otherwise, each experiment is repeated five times and each figure plots averages across all repeats. Error bars are used to indicate the Standard Error of Measurement (SEM).
+Here, the Transformer learns the deterministic breaking rules implicitly encoded in the data.
+Evaluating such a model is simple, because the generated sequences need to exactly match the sequences generated by $A^{*}$.
+  2. 2.
+    Non-deterministic $A^{*}$ dataset: Sequences are generated by executing $A^{*}$ in a non-deterministic fashion (by randomly ordering child nodes and breaking equal cost ties randomly).
+Consequently, given a task prompt, the optimal plan and $A^{*}$ execution trace is no longer unique and there are multiple correct responses.
+Here, the Transformer learns to generate the random tie breaking rules implicitly encoded in the sequence data.
+Consequently, the generated sequences vary between different executions, but the resulting plans are still optimal and execution traces still respect $A^{*}$’s cost and heuristic calculations as described in Section 3.3.
+Figure 7 in Appendix C presents an overview of the token sequence length for each dataset and shows that the generated $A^{*}$ execution traces grow in length with task complexity.
+Figure 8 shows that training and test sets are matched in difficulty and have comparable trace lengths.
+For each task, one model may generate a search sequence ending either in an optimal plan, a feasible plan (a plan that is correct but sub-optimal), or an invalid plan.
+In Appendix D we outline how each model’s ability to predict a feasible and optimal plan is scored and details about how the search dynamics of the search-augmented and Searchformer models is evaluated.
+Unless indicated otherwise, each experiment is repeated five times and each figure plots averages across all repeats. All reported errors indicate the Standard Error of Measurement (SEM).
 4.1 Maze navigation
-(a) Deterministic Prediction
-(b) Non-deterministic Prediction
+In the first experiment set, we train a set of encoder-decoder Transformer models to predict optimal plans for maze navigation tasks.
+We vary the training dataset size and model size (the number of optimized parameters) between different training runs and evaluate each model on the test tasks generated using the same hyper-parameters.
+(a) Deterministic case
+(b) Non-deterministic case
+(c) Performance across task difficulties
 Figure 2: Predicting intermediate computation steps leads to robust performance in the low data regime.
-Each model was trained to predict optimal plans for $30\times 30$ maze navigation tasks.
-The decoder of the search-augmented models were trained on sequences of the format <trace><plan> whereas the decoder of the solution-only models were trained on sequence of the format <plan>.
-For each model, the number of free parameters (indicated in millions of parameters with “15M”, “46M”, and “175M”) is varied by changing the hyper-parameters of the encoder-decoder architecture.
-Appendix A.2 describes the used architecture and hyper-parameters in more detail.
-In the first experiment set, we train a set of encoder-decoder Transformer models to predict optimal plans for $30\times 30$ mazes.
-We vary the training dataset size and model size (the number of optimized parameters) between different training runs and evaluate each model on the test tasks generated using the same hyper-parameters (e.g., maze size).
-Evaluation with the exact-match criterion
+For each model, the number of free parameters (indicated in millions of parameters with “15M”, “46M”, and “175M”) is varied.
+2(a): Comparison of how many test tasks were answered with a correct token sequence when training on the deterministic $A^{*}$ dataset (exact-match criterion in Appendix D).
+2(b): Comparison for how many test task at least one optimal plan was found when training on the non-deterministic $A^{*}$ dataset (any-optimal-64 criterion in Appendix D).
+2(c): Performance degradation when increasing task difficulty (maze size).
+Here, the non-deterministic $A^{*}$ dataset was used and models are evaluated with the any-optimal-64 criterion.
+Deterministic $A^{*}$
 Figure 2(a) plots for how many test tasks a correct response was generated.
-Both solution-only and search-augmented models are trained on the deterministic $A^{*}$ dataset and follow the exact-match criterion.
+Both solution-only and search-augmented models are trained on the deterministic $A^{*}$ dataset and are evaluated if they exactly re-produce the token sequences generated with $A^{*}$ search (please refer to the exact-match criterion in Appendix D).
 One can observe that the solution-only model is outperformed by most search-augmented models.
@@
 In the low training data regime (100,000 training sequences and less), performance of the solution-only model degrades significantly, while the performance of each search-augmented model stays relatively high.
-This result is surprising, because the search-augmented models must generate a sequence of the format <trace><plan>, which is on average more than 10 times longer than the corresponding sequence of the format <plan> generated by the solution-only models.
-Furthermore, even the smallest search-augmented model with 15 million (15M) parameters significantly outperforms the much larger solution-only model with 175 million (175M) parameters.
-Evaluation with the any-optimal-64 criterion
-When trained on non-deterministic $A^{*}$ data, there could be multiple different optimal paths for one task, and in practice one would only use the model to compute an plan, the any-optimal-64 metric is more appropriate.
-Note that using the any-optimal-64 criterion results in higher absolute numbers compared to exact-match, because to score one point under this criterion only one generated response needs to be correct.
-Figure 5: Solution-only model performance.
-Performance of the solution-only models is primarily influenced by the number of training sequences.
-Increasing a model’s size does not always improve performance.
-Figure 6: Performance across different task difficulties.
-Increasing task difficulty from left to right degrades performance of the solution-only models significantly.
-Here, each model was trained on 50,000 sequences.
-Figure 4(b) plots the any-optimal-64 criterion on the test set as a function of the number of non-deterministic $A^{*}$ training sequences.
-We can observe a similar pattern: even the smallest search-augmented models outperform solution-only model, especially for small training set.
+This result is surprising, because for more than 90% of the test mazes, the search-augmented models generate <trace><plan>-formatted sequences that are thousands of tokens long without predicting any single token incorrectly.
+Moreover, the solution-only models, that on average predict ten times shorter sequences, are significantly outperformed by the search-augmented models.
+Even the smallest search-augmented model significantly outperforms the much larger solution-only model parameters.
+This result highlights the power of training Transformers to generate long algorithm execution traces.
+We do not observe compounding prediction errors that usually limit deep model-based RL agents (Asadi et al., 2018), because the used backward-causal decoder network constructs for an $n$-step sequence an $n\times n$ attention map.
+Here, this property of the Transformer architecture is used to boost performance when predicting an optimal plan.
+Non-deterministic $A^{*}$
+When trained on non-deterministic $A^{*}$ data, the model could output multiple different optimal paths for one task.
+Here, we use each model to generate 64 token sequences for each task.
+The test task is counted as correctly answered of any one of the 64 sequences contains an optimal plan (please refer to the any-optimal-64 criterion in Appendix D).
+Because we only test if at least one generated sequence contains an optimal plan, we obtain higher absolute numbers in Figure 2(b) than in Figure 2(a).
+Figure 2(b) plots for how many test tasks an optimal plan was found when generating for each test task 64 token sequences.
+Here, we can observe a pattern similar to the deterministic $A^{*}$ dataset: even the smallest search-augmented models outperform solution-only model, especially for a small training set.
 Moreover, we found that model size only impacts the performance of each of the search-augmented models when using very small training datasets (50,000 training sequences).
 For larger training dataset sizes no significant difference is found.
-Increasing the number of parameters of the solution-only models does not significantly improve their performance in the low data regime (Figure 6).
+Increasing the number of parameters of the solution-only models does not significantly improve their performance in the low data regime (Figure 9 in Appendix F).
 Performance across different task difficulty levels
-Figure 6 illustrates how a task’s difficulty influences the performance of each model.
+Lastly, Figure 2(c) illustrates how a task’s difficulty influences the performance of each model.
 Here, we focus again on the dataset generated by non-deterministic $A^{*}$, and consider the number of correctly solved test tasks as a function of maze size.
 The larger the maze, the larger the task’s state space and the more computation is required to find an optimal solution plan.
-While the solution-only model reaches close to the performance of search-augmented model for the small $10\times 10$ mazes, its performance drops rapidly as the task becomes more challenging.
-In contrast, the search-augmented models maintain their high accuracy as the task becomes harder, even for its smallest model size.
-Appendix A.5 presents A full comparison across all maze sizes.
-Figure 7: Performance comparison with the Sokoban dataset.
-Percentage of correctly answered test prompts for each model size.
-All models see 100,000 training sequences.
-The plot averages percentage values across three different seeds and evaluates each model on the same 200 test tasks.
-Test tasks are not included in the training dataset.
+While the solution-only model’s performance drops rapidly as the tasks become more challenging, the search-augmented models maintain a comparably high accuracy, even for its smallest model size.
+Appendix F presents a full comparison across all maze sizes.
 Overall, while the solution-only models learn to predict an optimal plan if the used training dataset is large and diverse enough, search-augmented models perform significantly better in the low data regime and scale better to more difficult tasks.
 The search-augmented models reach higher performance because they can perform on-demand computation during inference.
-More specifically, the search-augmented models can control the search dynamics for a grounded reasoning chain that leads to an optimal plan, while the solution-only models have to infer direct correlations between a task description and its optimal plan through supervised learning where many of such correlations can be spurious and unreliable in the test time.
+More specifically, the search-augmented models imitate the search dynamics for a grounded reasoning chain that leads to an optimal plan, while the solution-only models have to infer direct correlations between a task description and its optimal plan through supervised learning where many of such correlations can be spurious and unreliable during evaluation on the test task set.
 4.2 Solving Sokoban puzzles
-To test if similar results can be obtained on a different and more complex task with a different tokenization pattern, we generate a planning dataset for Sokoban puzzles.
-Here, $A^{*}$’s execution traces are generated non-deterministically and expressed as a token sequence as outlined in Figure 10.
-Figure 7 plots how often each model generated a correct optimal plan for each test task.
+Table 1: Test set performance in the Sokoban tasks.
+Over 200 unseen test Sokoban tasks, we report percentage of solved and optimally solved test tasks.
+For sequences ending in either an optimal and correct plan we report the SWC (Success Weighted by Cost) score, and ILR (Improved Length Ratio of Search Dynamics) scores.
+The better trace and solution quality, the higher the scores.
+Please check Appendix D for detailed definitions of these scores.
+Params. Model Solved (%) Optimal (%) SWC ($\uparrow$) ILR (solved, $\uparrow$) ILR (optimal, $\uparrow$)
+45M Solution only 90.3 ±$\pm 1.0$ 86.8 ±$\pm 0.3$ 0.890 ±$\pm 0.009$
+Search augmented 92.5 ±$\pm 1.0$ 90.8 ±$\pm 1.6$ 0.924 ±$\pm 0.011$ 0.908 ±$\pm 0.020$ 0.919 ±$\pm 0.019$
+Searchformer, step 1 95.5 ±$\pm 1.0$ 93.5 ±$\pm 1.0$ 0.953 ±$\pm 0.010$ 1.054 ±$\pm 0.025$ 1.062 ±$\pm 0.015$
+Searchformer, step 2 96.0 ±$\pm 0.5$ 93.4 ±$\pm 0.6$ 0.957 ±$\pm 0.005$ 1.158 ±$\pm 0.025$ 1.181 ±$\pm 0.012$
+Searchformer, step 3 95.5 ±$\pm 0.8$ 93.7 ±$\pm 1.6$ 0.953 ±$\pm 0.009$ 1.292 ±$\pm 0.044$ 1.343 ±$\pm 0.067$
+175M Solution only 95.7 ±$\pm 0.2$ 90.0 ±$\pm 0.8$ 0.949 ±$\pm 0.003$
+Search augmented 95.2 ±$\pm 0.9$ 93.2 ±$\pm 1.0$ 0.949 ±$\pm 0.010$ 0.925 ±$\pm 0.010$ 0.933 ±$\pm 0.011$
+757M Solution only 96.5 ±$\pm 0.1$ 92.2 ±$\pm 1.2$ 0.958 ±$\pm 0.002$
+To test if similar results can be obtained on a different and more complex task with a different tokenization pattern and different transition structure, we repeat our experiments for Sokoban puzzles using our non-deterministic $A^{*}$ implementation.
+Table 1 lists how often each model generated a correct optimal plan for each test task.
 As before, by training on execution traces, the search-augmented models outperform the solution-only models.
@@
 On average, this 747 million parameter solution-only model is still outperformed slightly by a smaller 175 million parameter search-augmented model.
-This experiment confirms the findings found for mazes on a dataset with more complex planning tasks that use a different tokenization method.
+This experiment further confirms our findings on more complex planning tasks with a different transition structure and a different tokenization method.
 4.3 Searchformer: Improving search dynamics via bootstrapping
-In this last experiment, we investigate how the search-augmented model can be iteratively improved to compute an optimal plan while predicting fewer search steps.
+In this last experiment, we investigate how the search-augmented model can be iteratively improved to compute an optimal plan while generating a shorter execution trace.
 Here, our goal is to shorten the length of the search trace while still producing an optimal solution.
-We start out with the smallest search-augmented model trained on the non-deterministic $A^{*}$ Sokoban dataset and use it to generate a new shorter sequence training dataset as outlined in Sec.
-3.3.
-For each Sokoban puzzle in the training data, we generated 32 different <trace><plan> sequences by sampling tokens from the Transformer’s output distribution and include the shortest generation (measured in tokens) if it contains an optimal plan.
+We start out with the smallest search-augmented model trained on the non-deterministic $A^{*}$ Sokoban dataset and use it to generate a new shorter sequence training dataset as outlined in Section 3.3.
+For each Sokoban puzzle in the training data, we generated 32 different <trace><plan>-formatted sequences by sampling tokens from the Transformer’s output distribution and include the shortest generation (measured in tokens) if it contains an optimal plan.
 Subsequently, we fine-tune the search-augmented model on this newly created training data (by running an additional 10,000 training steps) to obtain the first Searchformer model.
-Using this Searchformer model, we subsequently generate another short sequence dataset and repeat the fine-tuning procedure.
-This newly generated dataset is then used to further fine-tune the previously obtained Searchformer model for another 10,000 training steps.
-In total, we repeat this improvement step three times.
-(a) Comparison of distribution of search dynamics length between $A^{*}$ search and our models.
-Each test task contributes one data point of averaged length on trials that lead to optimal plans (i.e., average-on-optimal length (Sec.
-4.1.2)).
+Using this Searchformer model, we subsequently generate another short sequence dataset and repeat the fine-tuning procedure to further improve the model.
+(a) Search length improvement.
 (b) Distribution of average-on-optimal length.
-(c) Distribution of lengths of individual sequences that lead to an optimal plan.
-Figure 8: Improvement of search dynamics length on optimal plans via bootstrapping in Sokoban.
-After three improvement steps, our Searchformer model reduced the generated search dynamics token length to an average of 3094—a 26.8% reduction in total from 4222, the average $A^{*}$ search length.
-With each improvement step, the Searchformer models generated shorter and shorter token sequences when computing an optimal plan.
-8(a): Comparison of search dynamics length (in terms of number of tokens) between $A^{*}$ search and our models (search-augmented in blue, Searchformer step 1-3 in orange), on the test subset in which our models yield optimal plans.
-Here, for each test task, we average the lengths of sequences that ends in an optimal plan, out of 64 trials (i.e., average-on-optimal length (Sec.
-4.1.2)).
-The box plots show the distribution of average-on-optimal lengths, in which the left boundary, mid solid line, right boundary represents the 25%, 50% and 75% quantiles.
-Dotted lines are means and whiskers show the range.
-8(b): Histograms of average-on-optimal lengths.
-8(c): Histograms of the length of individual sequences that lead to optimal plan.
-Because it is more likely for each model to generate an optimal plan for easy tasks with fewer search steps, the distributions are skewed towards the left—towards shorter sequence lengths.
-The data plotted in this figure aggregates sequence lengths across three experiment repeats.
-Figure 8 illustrates how the sequence lengths generated by the Searchformer model’s are iteratively shortened by our bootstrapping method.
-With every improvement step, the length of the generated traces—the number of search steps—decreases (Figure 8(a)).
+Figure 3: Improvement of search dynamics length via bootstrapping in Sokoban 3(a): For each Sokoban test task that was answered with an optimal plan, the average generated execution trace length is computed and averaged.
+The $A^{*}$ reference values are computed by generating an equal number of execution traces for each test task.
+For each iteration, we compare the subset of test tasks that were answered with an optimal plan by the search-augmented and Searchformer model and compare the relative improvement over our $A^{*}$ reference implementation.
+Figure 13 (Appendix F) shows a box plot with more details about the sequence length distribution.
+3(b): Distribution of execution trace lengths for each optimally answered test task.
+Here, the sequence lengths were first averaged per test task and the resulting distribution is plotted in each histogram.
+The search-augmented model is trained to imitate our $A^{*}$ implementation.
+After three iterations of search dynamics bootstrapping, the resulting Searchformer model generates on average shorter execution traces.
+Figure 3(a) illustrates how the sequence lengths generated by the Searchformer model’s are iteratively shortened with our search dynamics boostrapping method.
+With every improvement step, the average length of the generated traces—the number of search steps—decreases (Figure 3(a)).
 When computing an optimal plan, the final Searchformer model generates search dynamics sequences that are on average 26.8% shorter than the sequences generated with $A^{*}$ search.
-Consequently, the Searchformer model found a more efficient way to plan a complicated task and go beyond $A^{*}$’s search, even for unseen test puzzles.
+Consequently, the Searchformer model found a way to find a plan in a complex task that is more efficient in terms of search steps than the $A^{*}$ implementation used to train the initial search-augmented model.
 In Figure 3(b) we can observe that the search-augmented model generates sequences that on average match the sequences generated with $A^{*}$ search in length.
 The Searchformer models generate shorter sequences resulting in a distribution that is skewed towards shorter sequence lengths.
-We find that on average each model produces an optimal plan more often for tasks that can be solved with fewer search steps.
-Generating an optimal plan for more complex tasks is possible with each model but requires the model to be prompted more often.
-This is illustrated in Figure 8(c) which plots the sequence length distribution for individual sequences.
-The sequence length distribution of both models is skewed towards shorter sequences and the models do not match the $A^{*}$ sequence length distribution.
-Table 1: Test set performance in the Sokoban tasks.
-Over 200 unseen test Sokoban tasks, we report percentage of solved, percentage of optimal solutions, Success weighted by cost (SWC) and Improved Length Ratio (ILR) w.r.t $A^{*}$ reference, averaged over sequences ending in an optimal/correct plan.
-Please check Sec.
-4.1 for more detailed definitions of optimality metrics (Sec.
-4.1.1) and search length metrics (Sec.
-4.1.2).
-The result shows that continue fine-tuning Searchformer yields shorter search dynamics length and at the same time improved percentage of solved/optimally solved task.
-All values are averages with errors indicating SEM.
-Params. Model Solved (%) Optimal (%) SWC ILR-on-solved ILR-on-optimal
-45M Solution only 90.3 $\pm 1.0$ 86.8 $\pm 0.3$ 0.890 $\pm 0.009$
-Search augmented 92.5 $\pm 1.0$ 90.8 $\pm 1.6$ 0.924 $\pm 0.011$ 0.908 $\pm 0.020$ 0.919 $\pm 0.019$
-Searchformer, step 1 95.5 $\pm 1.0$ 93.5 $\pm 1.0$ 0.953 $\pm 0.010$ 1.054 $\pm 0.025$ 1.062 $\pm 0.015$
-Searchformer, step 2 96.0 $\pm 0.5$ 93.4 $\pm 0.6$ 0.957 $\pm 0.005$ 1.158 $\pm 0.025$ 1.181 $\pm 0.012$
-Searchformer, step 3 95.5 $\pm 0.8$ 93.7 $\pm 1.6$ 0.953 $\pm 0.009$ 1.292 $\pm 0.044$ 1.343 $\pm 0.067$
-175M Solution only 95.7 $\pm 0.2$ 90.0 $\pm 0.8$ 0.949 $\pm 0.003$
-Search augmented 95.2 $\pm 0.9$ 93.2 $\pm 1.0$ 0.949 $\pm 0.010$ 0.925 $\pm 0.010$ 0.933 $\pm 0.011$
-757M Solution only 96.5 $\pm 0.1$ 92.2 $\pm 1.2$ 0.958 $\pm 0.002$
 As reported in Table 1, fine-tuning the model resulted in a significant performance improvement, reducing the rate of incorrect and non-optimal solutions by 40% and 30% respectively.
-After three steps of fine-tuning, the 45M Searchformer shows similar to a 757M model that learns to predict a plan (task solution) only.
-The comparison using the more fine-grained SWC measures, which also take sub-optimal plans into consideration, shows a similar trend.
-See Sec.
-4.1 for definitions of evaluation measures.
-Reduced length of search dynamics with Searchformer
-Lastly, we evaluate to which extend the search-augmented and Searchformer models reduced the number of generated search steps to compute a correct plan.
-This analysis is different from the results presented in Figure 8 because here we focus on computing length ratios between traces generated with each model and traces generated with $A^{*}$ search using Improved Length Ratio (ILR) measure.
-This improvement can be observed in Table 1 for each Searchformer model where with every improvement step the model generates shorter and shorter execution traces leading to higher ILR scores.
-Furthermore, the scores for the two search-augmented models are slightly below one, indicating that the search-augmented models tend to perform slightly more search steps than $A^{*}$ search, while fine-tuned models has a ILR score higher than 1, showing more efficient search dynamics, in particular for those that lead to optimal plans.
-For example, $A^{*}$ search dynamics is $\sim 34.3\%$ longer than that produced by Searchformer after 3 steps of fine-tuning.
-More details about model’s performance characteristics are presented in Appendix A.4.
-5 Conclusion, limitations, and future work
-Prior work (Momennejad et al., 2023; Valmeekam et al., 2023a) has found that LLMs, which are based on Transformer models, struggle with solving complex decision making tasks.
+The Success Weighted by Cost (SWC) score (Wu et al., 2019) factors in how many test tasks were solved correctly and how close the predicted plans are to the optimal length (Appendix D).
+Here, a perfect score would be one, and one can see in Table 1 that the comparably small Searchformer matches the performance of the largest solution-only model (also note the small SEM values).
+Furthermore, the Improved Length Ratio of Search Dynamics (ILR) measures how much the length of each execution trace is shortened (Appendix D).
+With each improvement iteration the scores increase and climb above one.
+For example, $A^{*}$ search dynamics is $\sim 34.3\%$ longer than the sequences generated by Searchformer after 3 steps of fine-tuning.
+The results reported in Figure 3 and in Table 1 only compare each model’s performance on test tasks that were either correctly or optimally solved.
+To test if a model overfits only on easier test tasks with shorter execution traces, we plot in Figure 10 (in Appendix E) the execution trace length generated with $A^{*}$ against the execution trace length generated by each model as a scatter plot.
+Each point in this plot corresponds to a single test task.
+Here, the trend of shortening the execution trace via search dynamics bootstrapping is clear and one can also observe that neither model only specializes on solving easier test tasks with shorter execution traces.
+5 Discussion
+Prior work (Momennejad et al., 2023; Valmeekam et al., 2023a) has found that LLMs struggle with solving complex decision making tasks.
 Searchformer demonstrates that with appropriate training data, Transformers can in fact solve complex planning tasks.
 Moreover, Searchformer robustly follows the intermediate steps—the execution trace—of a symbolic planner and improves (in terms of trace length) beyond the human-crafted rule-based planning strategy it was initially trained on.
-Compared to solution-only models that directly predict a solution, our search-augmented models require fewer training sequences and scale better to more difficulty tasks.
-Limitations
+Compared to solution-only models that directly predict a solution, our search-augmented models require fewer training sequences and scale better to more complex planning tasks.
+5.1 Limitations
 Currently, Searchformer is trained on the execution traces of $A^{*}$ to learn a complex planning strategy.
-However, the trace length may grow exponentially in the length of an optimal plan (see Figure 3), and training on the resulting token sequence data become computationally very expensive.
+However, the trace length may grow exponentially in the length of an optimal plan (see Figure 7), and training on the resulting token sequence data can become computationally very expensive.
 In fact, the presented experiments use token sequences that are significantly longer than the sequences used to train LLMs such as Llama 2 (Touvron et al., 2023).
-[^3] The 175 million parameter model trained on the Sokoban dataset uses 64 GPUs with 32GB of memory.
-Future work
-One way to mitigate this limitation is to use curriculum learning: starting from simple tasks with reasonably long execution traces, train and fine-tune the Searchformer to reduce the trace length, and then adapt the improved model to more complex tasks.
-Another possibility is to integrate better heuristics or value functions into $A^{*}$ search, similar to Monte Carlo Tree Search (MCTS) (Coulom, 2006), to cap the maximal depth the search algorithm explores.
+5.2 Future work
+One way to mitigate this limitation and improve the efficiency of the presented methods is to use curriculum learning: starting from simple tasks with reasonably long execution traces, train and fine-tune the Searchformer to reduce the trace length, and then adapt the improved model to more complex tasks.
+Another possibility is to explore other planning algorithms or integrate better heuristics or value functions into $A^{*}$ search, similar to MCTS, to cap the maximal depth the search algorithm explores.
 Integrating hierarchical planning methods and temporal abstractions (Sutton et al., 2023, 1999; Dietterich, 2000; Hafner et al., 2022) are another avenue.
 This would equip the resulting model with the ability to abstract over multiple time steps and states to find an optimal plan using fewer computation steps.
-We hope that our study informs further research about how to build Transformer-based models that can robustly solve decision making tasks and mitigate the limitations previously outlined by Momennejad et al.
-and Valmeekam et al..
+In comparison to Plansformer (Pallagani et al., 2022), the presented work demonstrates how to train Transformers from scratch to solve complex planning tasks on synthetic datasets.
+We believe that our results and methods could be combined with methods such as Plansformer to fine-tune LLMs and enable them to solve complex planning tasks more robustly.
+Ultimately, we hope that our study sheds light on how Transformers can be used for multi-step planning and we hope to inform further research about better understanding the reasoning capabilities of LLMs.
 6 Broader Impact
@@
 While the tasks we explored in this paper can be easily solved with simple symbolic solvers, it is important to study the effectiveness of neural networks on such tasks.
-Neural network based Large Language Models have shown impressive performance on non-symbolic tasks, such as following complex natural language instructions, but still fail on relatively simple symbolic tasks like adding numbers.
-While the results in this paper are encouraging, they are still far from being competitive with symbolic solvers.
+Here, we provide a proof of concept on how Transformer-based neural networks can be used to robustly solve complex planning tasks.
+With our work, we hope to inform further research into better understanding the reasoning capabilities of Large Language Models.
 Acknowledgment
-We would like to thank Qinqing Zheng for helpful discussions and comments on earlier versions of this paper. We would also like to thank Amy Zhang for helpful discussions on this work.
-Appendix A Appendix
-A.1 Training encoder-decoder Transformers to predict optimal plans
+We would also like to thank Amy Zhang for helpful discussions on this work.
+Appendix A Training encoder-decoder Transformers to predict optimal plans
 For this work, we consider encoder-decoder Transformers (Raffel et al., 2020) and process each prompt—the task specification—with an encoder network to obtain an embedding of a planning task.
@@
 We refer to a model predicting sequences of the format <trace><plan> as a search-augmented model and a model predicting sequences of the format <plan> as solution-only models.
-A.1.1 Encoder-decoder architecture
+A.1 Encoder-decoder architecture
 A token sequence is processed by a neural network by first indexing the set of all tokens contained in a specific dataset.
@@
 $g_{\boldsymbol{\theta}_{\text{dec}}}:\boldsymbol{z}_{1:n},y_{1:m}\mapsto\boldsymbol{p}_{1:m}.$ (2)
-The encoder network $f_{\boldsymbol{\theta}_{\text{enc}}}$ and decoder network $g_{\boldsymbol{\theta}_{\text{dec}}}$ internally both use a number of stacked causal attention layers to form an encoder-decoder Transformer (Vaswani et al., 2017) as outlined in more detail in Appendix A.2.
+The encoder network $f_{\boldsymbol{\theta}_{\text{enc}}}$ and decoder network $g_{\boldsymbol{\theta}_{\text{dec}}}$ internally both use a number of stacked causal attention layers to form an encoder-decoder Transformer (Vaswani et al., 2017) as outlined in more detail in Appendix B.
 We denote a concatenation of all encoder parameters $\boldsymbol{\theta}_{\text{enc}}$ and decoder parameters $\boldsymbol{\theta}_{\text{dec}}$ with the vector $\boldsymbol{\theta}$.
-A.1.2 Training with teacher forcing
+A.2 Training with teacher forcing
 An encoder-decoder architecture is optimized to generate responses that follow the distribution of a training dataset by minimizing the cross-entropy between the training data distribution $p_{\mathcal{D}}$ over prompt-response pairs $(x_{n},y_{m})$ and the distribution $p_{\boldsymbol{\theta}}$ with which the encoder-decoder model is generating responses.
@@
 This loss objective is optimized using gradient descent (Goodfellow et al., 2016, Chapter 10).
-A.2 Network architecture and hyper-parameters
 Figure 4: Attention blocks architecture used in the encoder-decoder Transformer architecture.
@@
 After the last decoder layer, hidden activations are mapped through a linear layer into a sequences of logits vectors to compute the next token probability vectors $\boldsymbol{p}_{1:m}$.
+Appendix B Network architecture and hyper-parameters
 The encoder-decoder Transformer first maps every token to a one-hot vector of the same dimension as the token vocabulary space.
@@
 All models used a RoPE frequency of 10000.
-Parameter 15M param. model 46M param. model 175M param. model 747M param. model
+Parameter 15M model 46M model 175M model 747M model
 Layers 6 8 9 16
@@
 Table 3: Optimization hyper-parameters.
-Every model was optimized using AdamW (Loshchilov and Hutter, 2019) with setting $\beta_{0}=0.9$ and $\beta_{1}=0.99$.
+Every model was optimized using AdamW (Loshchilov & Hutter, 2019) with setting $\beta_{0}=0.9$ and $\beta_{1}=0.99$.
 Initially, the learning rate was linearly interpolated: Starting at zero and then increasing linearly to the value listed below until step 2000.
-Then a cosine learning rate schedule was followed (Loshchilov and Hutter, 2016).
+Then a cosine learning rate schedule was followed (Loshchilov & Hutter, 2016).
 to the between zero at the first training step and the listed value below for the
@@
 All experiments were implemented in PyTorch 2.0 (Paszke et al., 2019) and default parameters were used unless reported here otherwise.
-A.3 Dataset generation
-bosworker 2 3box 2 4box 3 4dock 1 3dock 4 4wall 0 0wall 0 1...wall 6 6eosPromptboscreate worker 2 3 box 2 4 box 3 4 c0 c3close worker 2 3 box 2 4 box 3 4 c0 c3...create worker 5 4 box 2 3 c10 c3close worker 2 1 c12 c0Trace(2583 tokens)plan 2 3plan 1 3plan 1 4plan 1 5plan 2 5plan 2 4plan 3 4plan 2 4plan 2 3plan 2 2plan 1 2plan 1 1plan 2 1eosPlan
-Figure 10: Token sequence example for Sokoban This figure lists the token sequence for the Sokoban level depicted in Figure 2.
-(a) Plan Sequence Length
-(b) $A^{*}$ Execution Trace Length
-Figure 11: Sequence length distribution for each dataset. The training and test sets are designed such that their sequence lengths match closely.
+Appendix C Dataset generation
 All datasets were generated by first randomly sampling a task and then executing $A^{*}$ to obtain an optimal plan.
@@
 If the sampled task is solvable by A*, then the task was admitted to the dataset.
+Figure 5: Example Sokoban puzzle [^2] This level image was composed using image icons from https://github.com/morenod/sokoban (accessed 2023-11-21)..
+The prompt, $A^{*}$ execution trace, and optimal plan for this task is illustrated in  Figure 6 in Appendix C.
+For Sokoban, we only use a non-deterministic $A^{*}$ implementation.
 Due to the large volume of generated data, all data was stored in and transformed with a MongoDB (MongoDB Inc., ) instance and map-reduce techniques.
@@
 The test dataset contains plans and traces that are of comparable length to the training data (Figure 8).
-A.4 Searchformer performance analysis
-(a) Comparison of all sequences ending with an optimal plan
-(b) Comparison of all sequences ending with a correct plan
-(c) Comparison of all sequences ending with a correct but sub-optimal plan
-Figure 12: Sequence length comparison between sequences generated with $A^{*}$ search and sequences generated with each model.
-Each dot in each scatter plot corresponds to one specific test task.
-On the $x$-axis, we plot the average token sequence length when $A^{*}$ search is used.
-On the $y$-axis, we plot the average token sequence length when each model is used.
-Error bars indicate standard deviations.
-Percentages indicate the fraction of the generated sequences that are included in each scatter plot.
-12(a): Sequence length comparison for all test prompts for which an optimal plan was generated using the corresponding model.
-12(b): Sequence length comparison for all test prompts for which a correct plan was generated using the corresponding model.
-This plot aggregates across sequences ending in an optimal plan and sequences ending in a correct but sub-optimal plan.
-12(c): Sequence length comparison for all test prompts for which a correct but sub-optimal plan was generated using the corresponding model.
-This plot only aggregates across sequences ending in a correct but sub-optimal plan.
-Sequences ending in an optimal plan were withheld.
-In Sec.
-3.3, each search-augmented and Searchformer model is evaluated by generating 64 token sequences for each Sokoban test task.
+bosworker 2 3box    2 4box    3 4dock   1 3dock   4 4wall   0 0wall   0 1...wall   6 6eosPromptboscreate worker 2 3 box 2 4 box 3 4 c0  c3close  worker 2 3 box 2 4 box 3 4 c0  c3...create worker 5 4 box 2 3         c10 c3close  worker 2 1                 c12 c0Trace(2583 tokens)plan 2 3plan 1 3plan 1 4plan 1 5plan 2 5plan 2 4plan 3 4plan 2 4plan 2 3plan 2 2plan 1 2plan 1 1plan 2 1eosPlan
+Figure 6: Token sequence example for Sokoban This figure lists the token sequence for the Sokoban level depicted in Figure 5.
+Figure 7: Training Sequence Length Comparison.
+The left panel plots the length of the solution-only sequences and the right panel plots the length of the search-augmented sequences, excluding the start and end of sequence tokens bos and eos.
+The whiskers indicate the range of all sequence lengths and the box plots indicate the 25%, 50%, and 75% quantiles.
+Because we focus on planning in complex sequential decision making tasks, the token sequences are multiple orders of magnitude longer than usual token sequences used to train LLMs—especially when $A^{*}$ execution traces are included in the responses.
+For example, fine-tuning of the Llama 2 model on human preference data is performed with sequences that are on average 600 tokens long (Touvron et al., 2023).
+(a) Plan Sequence Length
+(b) $A^{*}$ Execution Trace Length
+Figure 8: Sequence length distribution for each dataset. The training and test sets are designed such that their sequence lengths match closely.
+Appendix D Evaluation criteria
+In the presented experiments we evaluate if each model outputs an optimal or feasible plan and how long the generated search sequences are for the search-augmented and Searchformer models.
+D.1 Measuring plan optimality
+We evaluate whether the output plan is optimal using one of three criteria:
+  • Exact-match criterion.
+For each task, if the generated sequence from a trained model matches the output of deterministic $A^{*}$ exactly, it is labelled as correct, otherwise labelled as incorrect.
+This is only used to evaluate supervised cloning of deterministic $A^{*}$.
+  • Any-optimal-64 criterion.
+For each task, we sample 64 responses from a trained model.
+Each response is parsed and evaluated if it contains a feasible or optimal plan, regardless of the generated <trace> part.
+If any of the 64 plans is feasible and optimal, then the task is labelled as correct.
+  • SWC score.
+To further measure the sub-optimalness of the resulting plans, we also report the Success Weighted by Cost (SWC) (Wu et al., 2019), a statistic that factors in how close the cost $l_{i}$ of the best predicted correct plan (over 64 trials) is to the optimal plan cost $l_{i}^{*}$, averaged across all $n$ test tasks and weighted by a binary variable $c_{i}\in\{0,1\}$:
+    $\text{SWC}:=\frac{1}{n}\sum_{i=1}^{n}c_{i}\frac{l_{i}^{*}}{\max\{l_{i},l_{i}^{*}\}}.$
+    When computing the SWC score, the binary variable $c_{i}$ is set to one if a correct plan was found and zero otherwise.
+This SWC score lies between zero and one.
+If all generated sequences end in an optimal plan, then this value is one.
+D.2 Measuring search dynamics length
+For sequences ending in an optimal or feasible plan, we evaluate the length of sequence dynamics in terms of number of tokens using one of the two criteria:
+  • Average-on-optimal length. For each task, we sample 64 responses from a trained model and compute averaged search dynamics length for sequences that lead to an optimal plan.
+  • ILR score.
+To further measure how much improvement of the model-generated search dynamics against $A^{*}$ planning, we report the Improved Length Ratio of Search Dynamics (ILR) score.
+Specifically, for each test task $i$, we compute the ratio between the length $t_{i}$ of the shortest generated search dynamics sequence and the $A^{*}$ token sequence length $t_{i}^{*}$.
+We then average across all test tasks while only including ratios for tasks for which either an optimal or a correct (and potentially sub-optimal) plan was found, as specified by the binary variable $c_{i}$.
+The corresponding measures are thus called ILR-on-optimal and ILR-on-solved.
+The ILR is defined as
+    $\text{ILR}:=\frac{1}{n}\sum_{i=1}^{n}c_{i}\frac{t_{i}^{*}}{t_{i}}.$
+    The ILR measure can take non-negative values and values above one indicate that the model generates shorter search dynamics than the $A^{*}$ reference.
+Consequently, if the numbers lie significantly above one, then the model has found a more efficient way to search a task’s state space to compute an optimal plan.
+Appendix E Searchformer performance analysis
+In Section 3.3, each search-augmented and Searchformer model is evaluated by generating 64 token sequences for each Sokoban test task.
 For the same test task the same model can generate sequences that end in an optimal plan, an incorrect plan, or a correct but sub-optimal plan.
@@
 Statistically, the differences between Figure 10(a) and Figure 10(b) are marginal.
-A.5 Supplemental maze experiments
+Appendix F Supplemental figures for experiments
+Figure 9: Solution-only model performance.
+Performance of the solution-only models is primarily influenced by the number of training sequences.
+Increasing a model’s size does not always improve performance.
+(a) Comparison of all sequences ending with an optimal plan
+(b) Comparison of all sequences ending with a correct plan
+(c) Comparison of all sequences ending with a correct but sub-optimal plan
+Figure 10: Sequence length comparison between sequences generated with $A^{*}$ search and sequences generated with each model.
+Each dot in each scatter plot corresponds to one specific test task.
+On the $x$-axis, we plot the average token sequence length when $A^{*}$ search is used.
+On the $y$-axis, we plot the average token sequence length when each model is used.
+Error bars indicate standard deviations.
+Percentages indicate the fraction of the generated sequences that are included in each scatter plot.
+10(a): Sequence length comparison for all test prompts for which an optimal plan was generated.
+10(b): Sequence length comparison for all test prompts for which a correct plan was generated.
+This plot aggregates across sequences ending in an optimal plan and sequences ending in a correct but sub-optimal plan.
+10(c): Sequence length comparison for all test prompts for which a correct but sub-optimal plan was generated using the corresponding model.
+This plot only aggregates across sequences ending in a correct but sub-optimal plan.
 (a) Deterministic Plan Prediction
@@
 Training an encoder-decoder Transformer to imitate A* execution traces significantly boosts performance.
-14(a): Percentage of correctly generated responses.
+12(a): Percentage of correctly generated responses.
 Note that the plan only models only need to generate a few hundred token long plan to answer a prompt correctly.
@@
 If one token is incorrectly predicted, the response is counted as incorrect.
-14(b): Percentage of correctly generated plans.
+12(b): Percentage of correctly generated plans.
 Each model was used to generate multiple responses for each prompt.
 If one of the generated responses contains an optimal plan, the test prompt is counted as correctly answered.
-Figure 13 and Figure 14 plot how often a correct answer was provided by each model for all studied maze shape, training dataset size, and model parameterization combinations.
+Figure 13: Comparison of search dynamics length (in terms of number of tokens) between $A^{*}$ search and our models (search-augmented in blue, Searchformer step 1-3 in orange), on the test subset in which our models yield optimal plans.
+Here, for each test task, we average the lengths of sequences that ends in an optimal plan, out of 64 trials (i.e., average-on-optimal length (Appendix D)).
+The box plots show the distribution of average-on-optimal lengths, in which the left boundary, mid solid line, right boundary represents the 25%, 50% and 75% quantiles.
+Dotted lines are means and whiskers show the range.